### v2

# FUSION: Forecast-Embedded Agent Scheduling with Service Incentive Optimization over Distributed Air-Ground Edge Networks

Abstract
In this paper, we introduce a first-of-its-kind forecasting-driven, incentive-inherent service provisioning framework for distributed air-ground integrated networks that explicitly accounts for human-machine coexistence. In our framework, vehicular-UAV agent pairs (APs) are proactively dispatched to overloaded hotspots to augment the computing capacity of edge servers (ESs), which in turn gives rise to a set of challenges that we jointly address: highly uncertain spatio-temporal workloads, spatio-temporal coupling between road traffic and UAV capacity, forecast-driven contracting risks, and heterogeneous quality-of-service (QoS) requirements of human users (HUs) and machine users (MUs). To address these challenges, we propose FUSION, a two-stage optimization framework, consisting of an offline stage and an online stage. In the offline stage, a liquid neural network-powered module performs multi-step spatio-temporal demand forecasting at distributed ESs, whose outputs are exploited by an enhanced ant colony optimization-based routing scheme and an auction-based incentive-compatible contracting mechanism, to jointly determine ES-AP contracts and pre-planned service routes. In the online stage, we formulate the congestion-aware task scheduling as a potential game among HUs, MUs, and heterogeneous ES/UAVs, and devise a potential-guided best-response dynamics algorithm that provably converges to a pure-strategy Nash equilibrium. Experiments on both synthetic and real-world datasets show that FUSION consistently achieves higher social welfare and improved resource utilization, while maintaining latency and energy costs comparable to state-of-the-art baselines and preserving individual rationality, budget balance, and near-truthfulness.
Index Terms:
Distributed edge computing, liquid neural network, auction, potential games, heterogeneous network users.
I Introduction
Recent years have witnessed a new generation of intelligent applications such as real-time distributed video analytics, intelligent transportation systems, and large-scale mobile crowdsensing, flourishing across smart cities [1, 2]. Although edge servers (ESs) can provide computing capabilities for these intelligent applications at the network edge, their limited local resources make them susceptible to resource shortages and task congestion during peak demand periods [3].
To address the limitations of ESs, low-altitude edge computing enabled by unmanned aerial vehicles (UAVs) has received significant attention [4, 5]. Unlike static ESs, UAVs provide high mobility and flexibility, allowing them to be rapidly dispatched to hotspot regions to provide the needed computing resources. In this work, we highlight an important yet often-overlooked aspect of low-altitude edge systems: most existing studies implicitly assume that UAVs can always fly directly to target regions[8, 7, 9, 6], without considering their deployment origins. This is despite the fact that long travel distances due to poorly planned deployment origins can easily deplete UAVs’ batteries before reaching target locations. To fill this gap, we propose an air-ground cooperative edge computing paradigm: during periods of intensive demand, a set of ground vehicles are used to transport UAVs (vehicles and UAVs are collectively referred to “agent pairs (APs)”) to the vicinities of overloaded ESs. UAVs are then launched from strategically selected locations to deliver temporary computational augmentation for ESs. Despite its promising potential, operationalizing this hybrid AP-ES deployment paradigm introduces several challenges, which must be addressed:
(A) Spatio-temporal demand uncertainty. Edge computing demands exhibit strong spatio-temporal variability, and the coexistence of human users (HUs) and machine users (MUs) further complicates the management of distributed workloads. This variability makes it difficult to determine when, where, and how UAV resources should be dispatched. In such networks, static provisioning or purely reactive scheduling often leads to mismatches between supply and demand, resulting in congestion at emerging hotspots and resource underutilization in low-demand regions [13, 12]. The key challenge therefore lies in accurately forecasting spatio-temporal demand patterns from historical requests and contextual information, so as to enable proactive UAV deployment.
(B) Spatio-temporal coupling of vehicular mobility and UAV deployment. Supporting geo-distributed ESs requires APs to rely on vehicular mobility to reach target locations, which is constrained by road topology and traffic conditions. Moreover, the number of UAVs that can be deployed by a vehicle is tightly coupled with its visiting sequence and dwell time at each ES, since UAV dispatching, service execution, and recovery jointly consume time and onboard resources. Consequently, vehicle–UAV joint deployment becomes a highly coupled optimization problem that must simultaneously account for travel costs, service time windows, UAV dispatching logistics, and spatially distributed service rewards.
(C) Predictive contracting and incentive-compatible trading. In the ES–AP hybrid network, APs operated by private entities must be compensated by ES operators through advance contracts that commit service capacity and pricing for future peak-demand periods before demand realization. However, forecasting errors and dynamic supply–demand variations introduce significant risks for both parties: ESs may suffer unmet demand, while APs may incur revenue loss due to resource misallocation. The central challenge is thus to design a reliable offline trading mechanism that is individually rational, incentive compatible, and budget feasible, while operating prior to demand realization and service execution [14].
(D) Online network coordination under human–machine coexistence. HUs and MUs compete for shared resources provided by local ESs and leased UAVs, yet exhibit fundamentally different QoS requirements. HUs are typically delay-sensitive due to interactive applications, whereas MUs emphasize timely task completion before deadlines. The resulting challenge is to design fair and efficient distributed resource scheduling mechanisms that jointly serve HUs and MUs while ensuring stable operation and provable convergence guarantees.
The above four challenges form the motivation of our work. Specifically, to address them, we develop a forecasting-driven agent scheduling and incentive optimization (FUSION) framework for distributed service provisioning. In FUSION, to address challenge (A), we design a proactive spatio-temporal demand prediction via liquid neural networks (Pro-LNN), which transforms historical service traces and contextual features into multi-step workload forecasts. Then, to address challenge (B), we propose an enhanced ant colony optimization for vehicular route planning (eACO-VRP) based on the predicted demand. Afterwards, to address challenge (C), we introduce offline auction-based incentive-compatible contracting (Off-AIC2) that matches ESs with proper APs and obtains payments and rewards. Finally, to address challenge (D), we formulate an efficient online potential game among HUs, MUs, and ES/UAV, and propose the potential game-based best-response dynamics (PG-BRD) method, which integrates potential-guided dual decomposition with asynchronous best-response dynamics to drive the system towards a pure-strategy Nash equilibrium (NE) under congestion and fairness constraints. To sum up, our major contributions can be summarized as follows:
  • We present one of the first air-ground integrated edge network models that jointly captures vehicular-UAV APs, ESs, and heterogeneous human-machine (i.e., HU-MU) computing demands within a single framework.
  • We formulate ES–AP–user service provisioning by accounting for spatio-temporal mobility constraints, stochastic workload evolution, and strategic agent behaviors. We unveil that the offline ES–AP contract design admits a combinatorial mixed-integer formulation, while the online interactions between service demanders (HUs and MUs) and providers (ESs and APs) leads to a finite ordinal potential game with a combinatorial action space. To tackle these formulations, we develop a two-stage solution consisting of a forecast-driven offline stage and a game-theoretic online stage.
  • For the offline stage, we develop a multi-step spatio-temporal demand forecasting framework via Pro-LNN with executable vehicle-UAV route planning through an enhanced ACO-based algorithm (i.e., eACO-VRP). To our knowledge, this represents one of the first applications of the emerging LNN paradigm to spatio-temporal workload forecasting for proactive decision-making in air-ground integrated networks.
  • To address forecast-driven risk and strategic interactions in the offline stage, we propose Off-AIC2, a novel auction-based offline ES-AP contracting mechanism that incorporates demand forecasts into demand-supply matching and resource pricing, with provable guarantees of individual rationality, incentive compatibility, and budget feasibility.
  • For the online stage, we model real-time task scheduling as a multi-player potential game involving HUs, MUs, and heterogeneous ES/UAV resources. We then design PG-BRD, a potential-guided best-response algorithm with convergence guarantees to a pure-strategy NE.
  • Through extensive evaluations, we show that FUSION consistently delivers higher social welfare while keeping the delay and energy costs on par with existing baselines.
II Related Work
II-A Service Provision in UAV-Assisted Edge Networks
There exist an extensive literature on improving the communication/computation performance and reducing the energy footprint of UAV-assisted edge computing systems [8, 6, 7, 9]. For instance, Wang et al. [6] designed a multi-agent deep reinforcement learning (DRL)-based resource scheduling scheme for UAV-assisted wireless edge networks; Xu et al. [7] studied minimizing the average inference delay over UAV-assisted computing systems via DRL-based resource allocation; Li et al. [8] investigated distributed/federated edge intelligence for UAV-assisted networks, while optimizing the UAV energy consumption and resource utilization; Zeng et al. [9] introduced an ES-assisted UAV navigation system that integrates DRL-based decision making with network-aware containerized resource allocation to maximize the quality of navigation.
Although the above studies offer valuable insights, they predominantly presume that UAVs are always nearby and can provide immediate service, often without explicitly modeling their underlying deployment units or origins. As a result, the joint treatment of the four challenges summarized in Sec. I remains largely unexplored, an effort undertaken in this work.
II-B Auction-Driven and Game-Theoretic Service Provisioning
Auctions have been widely used for resource provisioning from service providers to users. For instance, Ren et al. [15] investigated truthful auctions for dependent task scheduling over dynamic edge networks; Xu et al. [16] designed combinatorial auction-based deep neural network (DNN) inference for heterogeneous edge-cloud networks to maximize social welfare; Li et al. [17] developed an online auction for generative content scheduling; Lu et al. [18] proposed auction-based distributed/federated learning to improve the distributed model training speed and energy consumption in edge networks. While the above studies provide valuable insights, they are mostly reactive (i.e., decisions are made after demand realization), and typically assume fixed resources and stable demand. In contrast, UAV-assisted edge computing involves dynamic resources and stringent delay and energy constraints, calling for spatio-temporally-aware auctions with minimal online interactions between service providers and end users while accounting for uncertainties in both resource availability and demand. This gap is addressed by our Off-AIC2 framework.
Beyond auctions, potential games have also been explored as distributed optimization strategies for service provisioning[12, 13]. For example, Ding et al. [12] investigated task processing across user–edge–cloud networks as potential games; Zhou et al. [13] modeled computation offloading as a potential game with latency–energy trade-offs; Hsieh et al. [19] proposed a knapsack potential game for task scheduling; Wu et al. [20] studied task scheduling for multi-cell edge–cloud networks through potential games with NE guarantees. Despite their notable contributions, the above works typically consider a single/homogeneous user type, overlooking the fundamentally different QoS requirements of HUs and MUs, as well as their interactions with heterogeneous ESs. Consequently, the effectiveness of potential games in UAV-assisted platforms that support human-machine coexistence, spatio-temporal task windows, and multi-party interactions remains largely unexplored, an effort addressed via our developed PG-BRD.
To further highlight our contributions, Table I compares our approach with representative existing studies.
III Overview and Modeling
III-A Overview of Our Methodology
We consider an air-ground integrated network, wherein nodes are partitioned into service demanders and service providers.
1) Service demanders (SDs): SDs are users that generate tasks such as model training/inference [^1] We assume that each SD is associated with one task that requires computing services from SPs in every future trading round.. To capture different QoS needs under human-machine coexistence, we further classify SDs into two categories:
  • HUs, denoted by $\bm{\mathcal{U}}^{\mathrm{H}}=\{u_{1}^{\mathrm{H}},\dots,u_{|\bm{\mathcal{U}}^{\mathrm{H}}|}^{\mathrm{H}}\}$ are sensitive to instantaneous service latency, as applications such as AR/VR, interactive gaming, real-time video conferencing, and intelligent driving are interactive and experience-driven, where even a slight change in latency or QoS can leave a pronounced impact on their perceived utility.
  • MUs, denoted by $\bm{\mathcal{U}}^{\mathrm{M}}=\{u_{1}^{\mathrm{M}},\dots,u_{|\bm{\mathcal{U}}^{\mathrm{M}}|}^{\mathrm{M}}\}$, correspond to autonomous devices such as robots, mapping vehicles, and sensing units. MUs are primarily concerned about whether their computation tasks can be completed before a certain deadline, and are therefore less impacted by instantaneous latency, as long as their overall task feasibility is preserved (e.g., a robot’s image-analysis task is considered feasible as long as the inference results are obtained before the deadline).
TABLE I: A summary of related studies
(ST: Spatio-Temporal, Pred.: Predictive, Het.: Heterogeneous)
Reference Network Characteristics Service Delivery Participant
Static Dynamic ST Offline Online Pred. Het. users Het. ESs
[6] $\surd$ $\surd$ $\surd$
[7] $\surd$ $\surd$ $\surd$
[8] $\surd$ $\surd$ $\surd$
[9] $\surd$ $\surd$
[19, 12, 20] $\surd$ $\surd$ $\surd$
[13] $\surd$ $\surd$ $\surd$
[15] $\surd$ $\surd$ $\surd$ $\surd$ $\surd$
[16, 17] $\surd$ $\surd$ $\surd$ $\surd$
[18] $\surd$ $\surd$ $\surd$
This Work $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$
2) Service providers (SPs): We consider two types of SPs:
  • ESs, denoted by $\bm{\mathcal{S}}=\{s_{1},\dots,s_{|\bm{\mathcal{S}}|}\}$, are geo-distributed across the map and provide low-latency services to nearby users. Each ES operates with a limited pool of orthogonal subcarriers and finite resource capacity, inherently constraining the number of users that can be concurrently served. During peak demand periods, hotspot ESs are particularly susceptible to resource saturation, which can trigger substantial queuing delays or even admission blocking.
  • APs, denoted by $\bm{\mathcal{V}}=\{v_{1},\dots,v_{|\bm{\mathcal{V}}|}\}$, are formed by vehicles carrying multiple on-board charging UAVs. Due to the limited power of UAVs, vehicles first traverse pre-planned routes to transport UAVs to designated service/ES locations. Upon arriving at their contracted ESs, UAVs are deployed to offer extra computing resources. This procedure incurs operational costs, including vehicle repositioning, UAV dispatching/launching procedures, flight energy consumption, and maintenance overhead.
Under this two-sided demand-supply structure, SDs compete for reliable computing services, while SPs collaboratively offer computing resources under their capacity and operational limitations. Nevertheless, in such a dynamic and distributed ecosystem, ESs face uncertain service demand: the temporal and spatial distribution of service request peaks across ESs is highly stochastic, and the composition ratio between HUs and MUs in the service requests also varies over time. To capture the time-varying behavior, a practical transaction is discretized into $T$ time-slots, with index $t$, i.e., $t\in\left\{0,1,2,...,T\right\}$. Further, to enable resource provisioning between SDs and SPs, we adopt a two-stage resource allocation framework that explicitly separates an offline stage from an online stage. From the ES perspective, the offline stage operates before the actual trading round, when only historical records and statistical demand patterns are available. At this stage (detailed in Sec. IV), the identities and tasks of HUs and MUs in the next trading round are unknown, making it impossible to pre-plan precise AP-ES matchings. Accordingly, in the offline stage of Fig. 1, we train Pro-LNN to forecast future service demand for each ES, and, based on these predictions, jointly design AP-to-ES visiting routes, UAV deployment plans, and auction-based contracts between ESs and APs (i.e., Off-AIC2).
Figure 1: Schematic of FUSION over an air–ground integrated network.
In contrast, the online stage starts (detailed in Sec. V) when real computing tasks arrive and the actual identities and locations of SDs become known. With full knowledge of active participants and their instantaneous task requirements, the system performs real-time task–resource scheduling. In this stage, we model the interactions among heterogeneous SPs and SDs as a potential game and construct an ordinal potential that captures the social welfare. Each SD then updates its task offloading decision and each SP adjusts its service admission in a fully distributed manner, where the resulting potential-guided best-response dynamics converge to a pure-strategy NE, yielding an online and congestion-aware provisioning of resources.
III-B Modeling of SDs, SPs, and Contracts
Hereafter, we provide detailed modeling of SPs and SDs.
III-B1 Modeling of SDs
We model HUs and MUs as follows:
(i) HUs: Each HU $u_{m}^{\mathrm{H}}\in\bm{\mathcal{U}}^{\mathrm{H}}$ is modeled by a 5-tuple $u_{m}^{\mathrm{H}}=\{e_{m}^{\mathrm{H,tx}},\,r_{m}^{\mathrm{H}},\,d_{m}^{\mathrm{H}},\,f_{m}^{\mathrm{H,loc}},\,\bm{C}_{m}^{\mathrm{H}}\}$, where $e_{m}^{\mathrm{H,tx}}$ is its uplink transmit power (Watts); $r_{m}^{\mathrm{H}}$ is its required computing workload (CPU cycles); $d_{m}^{\mathrm{H}}$ is its input data size (bits); $f_{m}^{\mathrm{H,loc}}$ is its local CPU frequency; and $\bm{C}_{m}^{\mathrm{H}}$ represents the set of reachable ESs within the communication coverage of $u_{m}^{\mathrm{H}}$.
(ii) MUs: Each MU $u_{n}^{\mathrm{M}}\in\bm{\mathcal{U}}^{\mathrm{M}}$ is represented by a 6-tuple $u_{n}^{\mathrm{M}}=\{\tau_{n}^{\mathrm{ddl}},\,e_{n}^{\mathrm{M,tx}},\,r_{n}^{\mathrm{M}},\,d_{n}^{\mathrm{M}},\,f_{n}^{\mathrm{M,loc}},\,\bm{C}_{n}^{\mathrm{M}}\}$, where $\tau_{n}^{\mathrm{ddl}}$ is its task deadline time; $e_{n}^{\mathrm{M,tx}}$ is its transmit power; $r_{n}^{\mathrm{M}}$ and $d_{n}^{\mathrm{M}}$ denote its workload and data size, respectively. Also, $f_{n}^{\mathrm{M,loc}}$ denotes its CPU capability, and $\bm{C}_{n}^{\mathrm{M}}$ captures SPs that can serve $u_{n}^{\mathrm{M}}$.
III-B2 Modeling of SPs
We model ESs and APs as follows:
(i) ESs: Each ES $s_{i}\in\bm{\mathcal{S}}$ is represented by a 6-tuple: $s_{i}=\{t_{i}^{\mathrm{E,b}},f_{i}^{\mathrm{E}},\,e_{i}^{\mathrm{E}},\,G_{i},\,l_{i}^{\mathrm{E}}\}$, where $t_{i}^{\mathrm{E,b}}$ denotes the time-slot at which $s_{i}$ begins to request resource supply to mitigate the shortage of edge resources within its own coverage area; $f_{i}^{\mathrm{E}}$ is its computing capacity (e.g., CPU cycles/s); $e_{i}^{\mathrm{E}}$ is its local power consumption (Watts); $l_{i}^{\mathrm{E}}=(x_{i}^{\mathrm{E}},y_{i}^{\mathrm{E}},h_{i}^{\mathrm{E}})$ is its 3D location. In addition, each ES $s_{i}$ is presumed to have a set of orthogonal subcarriers that can serve at most $G_{i}$ SDs simultaneously.
(ii) APs: each AP $v_{k}\in\bm{\mathcal{V}}$ is modeled by a 7-tuple: $v_{k}=\{C_{k}^{\mathrm{A}},\,f_{k}^{\mathrm{A}},\,e_{k}^{\mathrm{comp}},\,e_{k}^{\mathrm{move}},\,e_{k}^{\mathrm{fly}},\,G_{k}^{\mathrm{A}},\,l_{k}^{\mathrm{A}}\}$, where $C_{k}^{\mathrm{A}}$ denotes the maximum number of UAVs carried by the vehicle; $f_{k}^{\mathrm{A}}$ captures the computing capability of each mounted UAV; $e_{k}^{\mathrm{move}}$, $e_{k}^{\mathrm{comp}}$, and $e_{k}^{\mathrm{fly}}$ denote the power consumption of ground movement, UAV computation, and UAV flight, respectively; $G_{k}^{\mathrm{A}}$ denotes the maximum number of SDs that each UAV of $v_{k}$ can serve through its limited orthogonal subcarriers; $l_{k}^{\mathrm{A}}\langle t\rangle{=}(x_{k}^{\mathrm{A}}\langle t\rangle,y_{k}^{\mathrm{A}}\langle t\rangle)$ is the ground location of the vehicle at time $t$.
III-B3 Modeling of Offline Contracts
In the offline stage, a trading contract is established between each ES $s_{i}$ and AP $v_{k}$, denoted as $\mathbb{C}^{\mathrm{E}\leftrightarrow\mathrm{A}}_{i,k}=\{N^{\mathrm{Trad}}_{i,k},p_{i,k}^{\mathrm{ES}},r_{i,k}^{\mathrm{ES}}\}$, where $N^{\mathrm{Trad}}_{i,k}$ represents the quantity of traded resources, $p_{i,k}^{\mathrm{ES}}$ indicates unit payment (per task) from $s_{i}$ to contractual AP $v_{k}$, and $r_{i,k}^{\mathrm{ES}}$ refers to the unit reward (per task) of $v_{k}$ for offering services.
IV Offline Stage: Forecasting-Driven Contracting and AP Deployment
We first delve into capturing the service leasing among different types of SPs, for which we systematically transform the handling of uncertain spatio-temporal service demand into proactive, incentive-compatible, and cost-aware AP deployment decisions. Our major efforts include: (i) a predictive workload forecasting module that estimates the spatiotemporal dynamics of demands across distributed ESs; (ii) a cost-aware vehicle routing strategy based on eACO, which constructs time-window-feasible candidate associations/matching of APs to ESs; (iii) a contract-design module that leverages an incentive-compatible auction mechanism to determine long-term cooperative agreements between ESs and APs, including winning trading pairs, allocated tasks, and corresponding payment rules.
We let $x^{\mathrm{Off}}_{i,k}$ denote the assignment/matching between ESs and APs, where $x^{\mathrm{Off}}_{i,k}=1$ indicates that ES $s_{i}$ successfully secures the service of AP $v_{k}$ and thereby signs an offline contract, and $x^{\mathrm{Off}}_{i,k}=0$ otherwise. For notational simplicity, we define $\bm{X}^{\mathrm{Off}}=\big\{x^{\mathrm{Off}}_{i,k}\mid s_{i}\in\bm{\mathcal{S}},v_{k}\in\bm{\mathcal{V}}\big\}$ as the assignment profile.
IV-A Utilities of ESs, AP, and Auctioneer
IV-A1 Utility of ESs
In the offline stage, as delivering edge services to users and leasing computing capacity can generate both costs and revenues for ESs, we model the utility of each ES $s_{i}$ as its net profit as follows:
$U^{\mathrm{ES,Off}}_{i}=\sum_{v_{k}\in\bm{\mathcal{V}}}x^{\mathrm{Off}}_{i,k}N_{i,k}^{\mathrm{Trad}}\left(\overline{p}_{i}-p^{\mathrm{ES}}_{i,k}\right),$ (1)
where $\overline{p}_{i}$ denotes the historical average payment earned by $s_{i}$ for providing edge services to users.
IV-A2 Utility of APs
We model the utility of each AP as a combination of four key components: (i) Vehicular moving cost: the energy cost incurred by the vehicle traveling to ES $s_{i}$ is given by: $c^{\mathrm{move}}_{i,k}=\tau^{\mathrm{move}}_{i,k}e^{\mathrm{move}}_{k}$, where $\tau^{\mathrm{move}}_{i,k}$ denotes the traveling time of the carrier vehicle when moving from its origin to the target ES $s_{i}$, which depends on the driving distance and the AP’s vehicular speed; (ii) UAV flying cost: the energy consumption during UAV flight is given by: $c^{\mathrm{fly}}_{i,k}=\tau^{\mathrm{fly}}_{i,k}e^{\mathrm{fly}}_{k}$, where $\tau^{\mathrm{fly}}_{i,k}$ denote the total UAV flight time spent shuttling between $s_{i}$ and its associated service region, which increases with the number and size of the tasks to be served; (iii) UAV computing cost: the energy consumed by each UAV while providing resources is given by: $c^{\mathrm{comp}}_{i,k}=\tau^{\mathrm{comp}}_{i,k}e^{\mathrm{comp}}_{k}=\frac{\overline{r}e_{i}^{\mathrm{E}}}{f_{k}^{\mathrm{A}}}$, where $\overline{r}$ denotes the historical average computational workload per task [^2] Since trading is conducted in the offline stage, APs do not know the exact sizes of the tasks they will eventually process and can only evaluate them based on historical average data. , and $\tau^{\mathrm{comp}}_{i,k}$ represents the effective computing time that the UAV dedicates to processing offloaded tasks, determined by the aggregate computational workload and the UAV’s computing capability; (iv) Income for service provisioning: the revenue obtained from contributing resources is denoted by $r^{\mathrm{ES}}_{i,k}$. The overall utility of $v_{k}$ can therefore be expressed as follows:
$U^{\mathrm{A,Off}}_{k}\hskip-2.84526pt=\hskip-2.84526pt\sum_{s_{i}\in\bm{\mathcal{{S}}}}x^{\mathrm{Off}}_{i,k}\omega_{1}\left(r^{\mathrm{{ES}}}_{i,k}\hskip-2.84526pt-c^{\mathrm{move}}_{i,k}\hskip-2.84526pt-c^{\mathrm{fly}}_{i,k}-N_{i,k}^{\mathrm{Trad}}c^{\mathrm{comp}}_{i,k}\right)+\omega_{2}c^{\mathrm{hard}},$ (2)
where $\omega_{1}$ and $\omega_{2}$ denote cost coefficients that convert energy consumption into monetary units (e.g., dollars), and $c^{\mathrm{hard}}$ represents the hardware depreciation/maintenance cost.
IV-A3 Utility of the Auctioneer
An auctioneer (e.g., a trusted third party), plays a pivotal role in coordinating interactions among APs and ESs, including bids/asks collection, determining winning APs-ESs associations, and trading contract design. We define the utility of auctioneer as the difference between the income obtained by APs and the expense paid by ESs:
$U^{\mathrm{Auc,Off}}=\sum_{s_{i}\in\bm{\mathcal{S}}}\sum_{v_{k}\in\bm{\mathcal{V}}}x^{\mathrm{Off}}_{i,k}N^{\mathrm{{Trad}}}_{i,k}\left(p^{\mathrm{{ES}}}_{i,k}-r_{i,k}^{\mathrm{ES}}\right).$ (3)
IV-B Problem of Interest in Offline Stage of FUSION
Our overarching goal during the offline stage of FUSION is maximizing the social welfare (SW) defined as the collective utilities of engaged parties (i.e., ESs, APs, auctioneer) as:
$U^{\mathrm{SW}}=\sum_{s_{i}\in\bm{\mathcal{S}}}U^{\mathrm{ES,Off}}_{i}+\sum_{v_{k}\in\bm{\mathcal{V}}}U^{\mathrm{A}}_{k}+U^{\mathrm{Auc}}.$ (4)
[…]
s.t. $\sum_{v_{k}\in\bm{\mathcal{V}}}x^{\mathrm{Off}}_{i,k}\leq 1,~~~\forall s_{i}\in\bm{\mathcal{S}}$ (5a)
$x^{\mathrm{Off}}_{i,k}\in\left\{0,1\right\},~~~\forall s_{i}\in\bm{\mathcal{S}},~\forall v_{k}\in\bm{\mathcal{V}}$ (5b)
$\sum_{s_{i}\in\bm{\mathcal{S}}}{x^{\mathrm{Off}}_{i,k}N^{\mathrm{{Trad}}}_{i,k}}\leq G_{k}^{\mathrm{A}},~~~\forall v_{k}\in\bm{\mathcal{V}}$ (5c)
[…]
IV-C Forecasting, Agent Deployment, and Contracting
We next present the offline stage that couples demand forecasting, AP deployment, and long-term contracting, with detailed interaction flow moved to Appx. B-A. In what follows, we describe the three modules, i.e., Pro-LNN-based demand forecasting, eACO-VRP-based AP deployment, and the Off-AIC2 contracting mechanism.
IV-C1 Design of Proactive Spatio-Temporal Demand Prediction via Liquid Neural Networks (Pro-LNN)
Unlike discrete-time recurrent neural networks (RNNs) or Transformers, LNNs [23] implement continuous-time dynamics with neuron-wise time constants and exponential-style state updates. This makes them particularly suitable for modeling time series that are sampled at irregular resource trading intervals, exhibit non-stationary spatio–temporal demand patterns, and must be processed under strict edge-computing resource constraints. In our setting, Pro-LNN takes as input the ES-side demand time series across trading rounds $\nu=1,2,\ldots$, where the time gaps $\Delta\nu$ between rounds can be irregular and some observations may be missing due to asynchronous reporting. The predictor therefore must stay stable under such irregular $\Delta\nu$ while capturing both short-term and long-term trends, motivating our LNN-based architecture.
Let $N_{i}^{\mathrm{Dem},(\nu)}$ denote the task demand of ES $s_{i}$ at the $\nu$-th trading round. We form an input vector $\mathbf{z}_{i}^{(\nu)}=\big[N_{i}^{\mathrm{Dem},(\nu)},\mathbf{x}_{i}^{(\nu)}\big]$, where $\mathbf{x}_{i}^{(\nu)}$ stacks optional exogenous features at round $\nu$, such as time-of-day and day-of-week indicators, ES-type and capacity descriptors, and local mobility/load statistics extracted from recent logs. Given a length-$L$ history of past observations $\{\mathbf{z}_{i}^{(\nu-L+1)},\ldots,\mathbf{z}_{i}^{(\nu)}\}$, Pro-LNN encodes this sequence via a liquid cell with neuron-wise time constants, integrates the measurements into a hidden state $\mathbf{h}_{i}^{(\nu)}$, and then applies a shallow readout network to produce an $H$-step forecast $\{\hat{N}_{i}^{\mathrm{Dem},(\nu+1)},\ldots,\hat{N}_{i}^{\mathrm{Dem},(\nu+H)}\}$. In a nutshell, Pro-LNN is trained with a composite loss, which is an $H$-step MSE between the forecasts and the ground-truth demands, and we additionally use a multi-step consistency regularizer that rolls the predictor forward in closed-loop over $H$ steps and penalizes the accumulated prediction errors at future trading rounds. These forecasts are computed offline before each trading round and then fed into the eACO-VRP module for ES recommendation and AP route pre-planning, as well as the auction module for ES–AP contract design. The detailed update equations and training pseudocode of our Pro-LNN are deferred to Appx. B-B to avoid detracting from the main discussion.
IV-C2 Enhanced Ant Colony Optimization for Vehicle Route Planning (eACO-VRP)
Within the proposed offline stage of FUSION, once Off-AIC2 has identified a feasible trading region and candidate ES–AP pairs, the auction platform still needs to decide, for each winning AP, which ESs should be physically visited and in what order, under time-window and capacity constraints. This induces a capacitated vehicle-routing problem with time windows and heterogeneous travel costs and utilities. Since our APs (e.g., UAVs or service vehicles) and the auction platform operate under tight energy and latency budgets and should periodically recompute routes as demand forecasts change, we adopt a lightweight ant colony optimization (ACO) paradigm and tailor it into an eACO-VRP that can be invoked as a subroutine inside Off-AIC2, balancing solution quality with low computational complexity and energy cost.
[…]
$pr^{\mathbbm{k}}_{r,s}=\begin{cases}\displaystyle\frac{\varphi^{\varepsilon_{1}}_{r,s}\,\eta^{\varepsilon_{2}}_{r,s}\,\big(1/\mathrm{width}_{s}\big)^{\varepsilon_{3}}}{\sum\limits_{q\in J^{\mathbbm{k}}(\mathrm{n}_{r})}\varphi^{\varepsilon_{1}}_{r,q}\,\eta^{\varepsilon_{2}}_{r,q}\,\big(1/\mathrm{width}_{q}\big)^{\varepsilon_{3}}}&\text{if }\mathrm{n}_{s}\in J^{\mathbbm{k}}(\mathrm{n}_{r}),\\[12.0pt] 0&\text{otherwise,}\end{cases}$
with positive weights $\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}$. In (6), $\varphi_{r,s}$ denotes the pheromone level on edge $(\mathrm{n}_{r},\mathrm{n}_{s})$ (see (7)); $\eta_{r,s}$ is the distance between $\mathrm{n}_{r}$ and $\mathrm{n}_{s}$. If the ant arrives early (i.e., before the time-slot when ES $s_{i}$ begins to request extra resources), it can pre-position the AP’s UAV at this ES; in this case, the remaining time $\mathrm{width}_{s}$ is defined as the difference between the arrival time and the time-slot when ES $s_{i}$ begins to request extra resources. Conversely, if the ant reaches $\mathrm{n}_{s}$ after the time-slot when ES $s_{i}$ begins to request extra resources, this ES is regarded as infeasible and eliminated in the route construction. In (6), $J^{\mathbbm{k}}(\mathrm{n}_{r})$ denotes the set of feasible ES vertices that ant $\mathbbm{k}$ can reach when standing at node $\mathrm{n}_{r}$: an ES $\mathrm{n}_{s}\in J^{\mathbbm{k}}(\mathrm{n}_{r})$ if it satisfies two feasibility conditions: (i) the current time plus the travel time from $\mathrm{n}_{r}$ to $\mathrm{n}_{s}$ should not exceed the time-slot when ES $s_{i}$ begins to request extra resources, and (ii) there must still be remaining UAVs at AP $v_{k}$ to deploy. Once a route is found by ants, pheromones on edges are updated as follows:
$\varphi_{r,s}\ \leftarrow\ (1-\theta)\,\varphi_{r,s}+\Delta\varphi_{r,s},$ (7)
where $0<\theta<1$ is the evaporation coefficient, and
$\Delta\varphi_{r,s}=\left\{\begin{aligned} &\sum_{(\mathrm{n}_{r},\mathrm{n}_{s})\in\bm{M}^{\mathrm{max}}}\,\Delta{U^{\mathrm{A}}_{r,s}},\quad&&\text{if }(\mathrm{n}_{r},\mathrm{n}_{s})\in\bm{M}^{\mathrm{max}},\\ &0,&&\text{otherwise},\end{aligned}\right$. (8)
captures the deposited pheromone, where $\triangle{U^{\mathrm{A}}_{r,s}}$ denotes the utility increment for $v_{k}$ when moving from $\mathrm{n}_{r}$ to serve the ES at $\mathrm{n}_{s}$, and $\bm{M}^{\mathrm{max}}=\Big\{\bm{M}^{\mathbbm{k}}\ \Big|\ \max~\sum_{(r,s)\in\bm{M}^{\mathbbm{k}}}\triangle{U^{\mathrm{A}}_{r,s}},\ \forall\,\mathbbm{k}\in\mathbb{K}\Big\}$ denotes the path with the largest utility among ants. For brevity, the pseudo-code of eACO-VRP is provided in Appx. B-C, while its four main steps are summarized below:
[…]
Step 2. Path exploration: At each iteration, an ant $\mathbbm{k}$ located at its current vertex $\mathrm{n}_{r}$ jointly evaluates the current time-slot, the travel time to each vertex, the service time at that vertex, and the time-slot when ES $s_{i}$ begins to request extra resources. This screening step yields the feasible candidate set $J^{\mathbbm{k}}(\mathrm{n}_{r})$ and the associated transition probabilities $pr^{\mathbbm{k}}_{r,s}$ as discussed above. The ant then probabilistically selects its next vertex using (6).
Step 3. Pheromone update: For each constructed path $\bm{M}^{\mathbbm{k}}$, we compute the cumulative utility $\sum_{(r,s)\in\bm{M}^{\mathbbm{k}}}\triangle{U^{\mathrm{A}}_{r,s}}$. Among all ants, the path with the largest utility is selected as $\bm{M}^{\mathrm{max}}$, and the pheromone trails on its edges are reinforced according to (7) and (8), while pheromones on other edges evaporate.
Step 4. Output: After finishing all the iterations, the route $\bm{M}^{\mathrm{max}}$ with the highest utility is returned as the pre-planned path for AP $v_{k}$, and the ES tasks located along this route are recommended as its preferred service targets.
IV-C3 Offline Auction-based Incentive-Compatible Contracting (Off-AIC2)
Building on the pre-planned AP routes obtained above, we now introduce Off-AIC2, an offline double-auction mechanism that establishes incentive-aware service contracts between ESs and APs. The goal is to decide ES–AP associations, how much service capacity should be reserved along each AP route, and how payments/rewards should be set so that participants (i.e., ESs and APs) have limited incentive to misreport their private values.
In Off-AIC2, on the buyer side, each ES $s_{i}$ predicts its total service demand $N_{i}^{\mathrm{Dem}}$ (e.g., the number of service units or tasks that need to be supported in the next trading round) and has a private per-unit valuation $v_{i}$ for receiving edge service. It then participates in Off-AIC2 by submitting a per-unit bid $\overline{\mathrm{bid}}_{i}$ together with its demand $N_{i}^{\mathrm{Dem}}$, where $\overline{\mathrm{bid}}_{i}$ represents the maximum payment it is willing to make per unit of service. On the seller side, each AP $v_{k}$, based on its maximum service capacity $C_{k}^{\mathrm{A}}$ (e.g., the total number of service units it can provide) and its service cost, reports a per-unit ask price $\mathrm{ask}_{k}$ and its capacity $C_{k}^{\mathrm{A}}$, where $\mathrm{ask}_{k}$ is the minimum acceptable payment per unit of service. Subsequently, for a potential ES–AP pair $(s_{i},v_{k})$, we define a pairwise utility $U_{i,k}^{\mathrm{Off}}=\mathrm{bid}_{i,k}-\mathrm{ask}_{k}$ ($U_{i,k}^{\mathrm{Off}}>0$ indicates a socially beneficial match). Off-AIC2 then uses the reported bids, asks, demands, and capacities to compute pairwise utilities and select a set of winning ESs and APs, determine their matching/association pattern, and compute payments and rewards. For brevity, the complete pseudo-code of Off-AIC2 is provided in Appx. B-D, while its main three steps are summarized below:
Step 1. List generation and pivotal-index search: Off-AIC2 operates as a one-shot sealed-bid double auction: all ES buyers $s_{i}$ and AP sellers $v_{k}$ simultaneously submit their bids, asks, and resource quantity information to an auctioneer without observing each others’ submissions. We first sort ESs into a list $\mathcal{L}^{\mathrm{ES}}$ in descending order of $\overline{\mathrm{bid}}_{i}$, and APs into a list $\mathcal{L}^{\mathrm{AP}}$ in ascending order of $\mathrm{ask}_{k}$. Without loss of generality, we relabel indices so that these sorted orders are directly reflected by $i$ and $k$. The algorithm then searches for a pair of pivotal ranks $(I^{\ast},K^{\ast})$ such that $\overline{\mathrm{bid}}_{I^{\ast}}\geq\mathrm{ask}_{K^{\ast}},\sum_{k=1}^{K^{\ast}}C_{k}^{\mathrm{A}}\geq\sum_{i=1}^{I^{\ast}}N_{i}^{\mathrm{Dem}}$.
In words, $(I^{\ast},K^{\ast})$ is the largest prefix of high-bidding ESs and low-asking APs whose aggregate demand and supply can be satisfied at a non-loss-making price level.
Step 2. ES–AP matching and route planning via eACO-VRP: Given the pivotal ranks $(I^{\ast},K^{\ast})$, the top-$I^{\ast}$ ESs and top-$K^{\ast}$ APs participate in the matching stage. We initialize the set of currently unmatched ESs as $\mathcal{U}^{\mathrm{free}}=\{1,\dots,I^{\ast}\}$ and iterate over AP sellers $v_{k}$ following the order in $\mathcal{L}^{\mathrm{AP}}$. For each AP $v_{k}$ with remaining capacity $C_{k}^{\mathrm{A}}>0$ and $\mathcal{U}^{\mathrm{free}}\neq\emptyset$, we construct a candidate ES set $\mathcal{C}_{k}=\big\{\,s_{i}\in\mathcal{U}^{\mathrm{free}}\,\big|\,U_{i,k}^{\mathrm{Off}}>0,\ \overline{\mathrm{bid}}_{i}\geq\mathrm{ask}_{k}\big\}$, where $U_{i,k}^{\mathrm{Off}}>0$ ensures that serving $s_{i}$ along some route of $v_{k}$ improves the system-level SW, and $\overline{\mathrm{bid}}_{i}\geq\mathrm{ask}_{k}$ guarantees that the ES’s per-unit value is larger than the AP’s ask price.
For such an AP $v_{k}$, we invoke the eACO-VRP routine (Sec. IV-C2), which considers ES demand volumes over time, travel time and energy cost of APs, and the payment that $v_{k}$ can receive, in order to jointly determine: (i) which ESs in $\mathcal{C}_{k}$ should be served without exceeding $v_{k}$’s remaining capacity $C_{k}^{\mathrm{A}}$, and (ii) in what visiting order these ESs are arranged along the route of $v_{k}$. By maximizing the AP’s utility over all such candidate routes, we obtain a selected ES set $\mathcal{K}_{k}^{\ast}\subseteq\mathcal{C}_{k}$ and the associated pre-planned path for AP $v_{k}$. Furthermore, for every $s_{i}\in\mathcal{K}_{k}^{\ast}$, we set $x_{i,k}^{\mathrm{Off}}=1$ and remove $s_{i}$ from $\mathcal{U}^{\mathrm{free}}$, ensuring that each ES is matched to (at most) one AP. Stacking these per-AP decisions over all $k$ yields the overall ES–AP assignment matrix $\bm{X}^{\ast}$. Moreover, since every successful pair $(s_{i},v_{k})$ satisfies $\overline{\mathrm{bid}}_{i}\geq\mathrm{ask}_{k}$, each winning ES is guaranteed not to pay more than its declared value in the matching stage, which contributes to ES-side individual rationality.
Step 3. Pricing and contract finalization via binary search: Once $\bm{X}^{\ast}$ is determined, Off-AIC2 computes ES payments and AP rewards through a monotone binary-search procedure that approximates critical-value pricing. The high-level goal is twofold: (i) discourage large deviations from truthful bidding, and (ii) keep the total ES payment and total AP reward closely aligned so that the mechanism does not incur a systematic deficit. In particular, for each matched ES buyer $s_{i}$ (with $\sum_{k}x^{\mathrm{Off}}_{i,k}=1$), we treat $\overline{\mathrm{bid}}_{i}$ as an effective bid parameter and search over the interval $\overline{\mathrm{bid}}_{I^{\ast}+1},\,\overline{\mathrm{bid}}_{i}]$ to find the smallest bid $p_{i}^{\mathrm{b}\ast}$ such that, if $s_{i}$ had declared this lower bid (while all other bids/asks remained unchanged) and Step 2 was re-executed, $s_{i}$ would still be selected as a winner. This $p_{i}^{\mathrm{b}\ast}$ is the ES’s critical payment: bidding any lower would cause $s_{i}$ to lose the contract. Symmetrically, for each matched AP seller $v_{k}$ (with $\sum_{i}x^{\mathrm{Off}}_{i,k}\geq 1$), we vary its effective ask within $\mathrm{ask}_{K^{\ast}+1},\,\mathrm{ask}_{k}]$ and re-run Step 2 to determine the largest reward level $r_{k}^{\mathrm{s}\ast}$ such that $v_{k}$ still remains in the winning set. This $r_{k}^{\mathrm{s}\ast}$ acts as a critical reward: if $v_{k}$ has a higher ask, it would not be matched to any ES.
The resulting settlement vectors $(\bm{p}^{\mathrm{b}\ast},\bm{r}^{\mathrm{s}\ast})$ finalize the offline contracts between ESs and APs under Off-AIC2. By tying payments and rewards to these critical values, the mechanism encourages near-truthful [^3] Unlike conventional auctions that ignore spatio-temporal aspects of the system, truthfulness in our setting is inherently shaped by spatial advantages: ESs close to AP routes can be profitably served with lower bids, whereas remote ESs incur higher travel and time costs and must bid higher to remain competitive; similarly, APs with route-aligned demand clusters face lower marginal costs than detour-intensive ones. As ES–AP matching is jointly coupled with route planning in eACO-VRP, strict dominant-strategy truthfulness cannot be guaranteed. Instead, the binary-search-based payment rule approximates critical-value pricing and induces near-truthful behavior in practice. reporting of private valuations while keeping buyer payments and seller rewards balanced at the system level (a property that we numerically demonstrate in Sec. VI-C3). In summary, Off-AIC2 satisfies individual rationality, approximate truthfulness, and budget balance. Since these properties follow standard definitions, we relegate their formal definitions and detailed proofs to Appx. B–E to maintain clarity and focus in the main text.
V Online Stage: Potential Games for Service Scheduling under Human-Machine Coexistence
Unlike the offline stage, the online stage aims to cope with real-time service requests generated by HUs and MUs. In practice, the number and composition of SDs that actually arrive in each trading round may deviate from the offline forecasts. As a result, the offline outcome can only reserve a feasible pool of UAVs and service capacities for ESs with heavy resource demands, but it cannot directly determine which SP should serve which SD, at what time, or at what service rate given the real-time state of the system. Therefore, in this section we build on the offline contracts and capacity budgets, and design an online service scheduling mechanism that dynamically refines SD–SP associations and resource allocation so as to adapt to the current system state.
In particular, we model the online service scheduling through interactions among two players: (i) SPs (e.g., ESs and APs), collected as $\mathcal{S}^{\mathrm{On}}=\{s^{\mathrm{On}}_{1},\ldots,s^{\mathrm{On}}_{\mathbbm{i}},\ldots,s^{\mathrm{On}}_{|\mathcal{S}^{\mathrm{On}}|}\}$; and (ii) SDs (e.g., HUs and MUs), collected as $\mathcal{U}^{\mathrm{On}}=\{u^{\mathrm{On}}_{1},\ldots,u^{\mathrm{On}}_{\mathbbm{j}},\ldots,u^{\mathrm{On}}_{|\mathcal{U}^{\mathrm{On}}|}\}$. In this stage, each SD strategically decides whether to execute its tasks locally or offload them to one of the available SPs, while SPs allocate their computing capacities to their associated tasks. Such mutually coupled decisions are then considered as a potential game. Specifically, to represent the offloading decision of $u^{\mathrm{On}}_{\mathbbm{j}}$, we define the following binary indicator vector:
$\hskip-8.53581pt\bm{X}^{\mathrm{On}}_{\mathbbm{j}}=\big(x^{\mathrm{On}}_{\mathbbm{j},0},\,x^{\mathrm{On}}_{\mathbbm{j},1},\,\ldots,\,x^{\mathrm{On}}_{\mathbbm{j},\mathbbm{i}},\,\ldots,\,x^{\mathrm{On}}_{\mathbbm{j},|\mathcal{S}^{\mathrm{On}}|}\big)\in\{0,1\}^{\,1+|\mathcal{S}^{\mathrm{On}}|},$ (9)
where $x^{\mathrm{On}}_{\mathbbm{j},0}=1$ indicates local task execution at $u^{\mathrm{On}}_{\mathbbm{j}}$ and $x^{\mathrm{On}}_{\mathbbm{j},\mathbbm{i}}=1$ indicates offloading to SP $s^{\mathrm{On}}_{\mathbbm{i}}$ (note that $\sum_{\mathbbm{i}=0}^{|\mathcal{S}^{\mathrm{On}}|}x^{\mathrm{On}}_{\mathbbm{j},\mathbbm{i}}=1,~\forall\,\mathbbm{j}$). Given the profile $\bm{X}^{\mathrm{On}}_{\mathbbm{j}}$, the resulting delays and costs are jointly determined by how many SDs concurrently choose the same SP and how its computing resources are shared. We next formalize this joint effect through a delay and congestion model, which will serve as the basis for defining the utilities of SDs and SPs.
V-A Utility of SDs and SPs
V-A1 Delay and Congestion Model
We consider the SP set $\mathcal{S}^{\mathrm{On}}$ and the SD set $\mathcal{U}^{\mathrm{On}}=\bm{\mathcal{U}}^{\mathrm{H}}\cup\bm{\mathcal{U}}^{\mathrm{M}}$, where $\bm{\mathcal{U}}^{\mathrm{H}}$ and $\bm{\mathcal{U}}^{\mathrm{M}}$ denote the sets of HUs and MUs, respectively, where each SP $s^{\mathrm{On}}_{\mathbbm{i}}$ has a finite computing capacity $f^{\mathrm{On}}_{\mathbbm{i}}>0$ (CPU cycles/s). When a subset of SDs $\mathcal{U}_{\mathbbm{i}}=\{u^{\mathrm{On}}_{\mathbbm{v}}\in\mathcal{U}^{\mathrm{On}}:x^{\mathrm{On}}_{\mathbbm{v},\mathbbm{i}}=1\}$ simultaneously offload their workload to $s^{\mathrm{On}}_{\mathbbm{i}}$, the SP must multiplex its limited computing resource among heterogeneous requests with different priorities and contract terms. To capture both congestion effects and differentiated service levels in a tractable way, we adopt a weighted processor sharing (WPS) rule [13]: each SD $u^{\mathrm{On}}_{\mathbbm{j}}$ associated with $s^{\mathrm{On}}_{\mathbbm{i}}$ is assigned a positive weight $\gamma_{\mathbbm{j},\mathbbm{i}}>0$ that reflects its service priority, contract tier, or fairness quota. The effective computing rate that $u^{\mathrm{On}}_{\mathbbm{j}}$ receives from $s^{\mathrm{On}}_{\mathbbm{i}}$ is then given by $\hat{f}_{\mathbbm{j},\mathbbm{i}}=f^{\mathrm{On}}_{\mathbbm{i}}\frac{\gamma_{\mathbbm{j},\mathbbm{i}}}{\sum_{u^{\mathrm{On}}_{\mathbbm{v}}\in\mathcal{U}^{\mathrm{On}}_{\mathbbm{i}}}\gamma_{\mathbbm{v},\mathbbm{i}}}$, and the corresponding edge-side computing delay can be obtained as:
$t^{\mathrm{comp}}_{\mathbbm{j},\mathbbm{i}}=\frac{r_{\mathbbm{j}}}{\hat{f}_{\mathbbm{j},\mathbbm{i}}}=\frac{r_{\mathbbm{j}}}{\gamma_{\mathbbm{j},\mathbbm{i}}}\times\frac{\sum_{u^{\mathrm{On}}_{\mathbbm{v}}\in\mathcal{U}^{\mathrm{On}}_{\mathbbm{i}}}\gamma_{\mathbbm{v},\mathbbm{i}}}{f^{\mathrm{On}}_{\mathbbm{i}}},$ (10)
where $r_{\mathbbm{j}}$ denotes the computing workload (e.g., CPU cycles needed) of $u^{\mathrm{On}}_{\mathbbm{j}}$. This WPS model compactly captures how the delay of each SD grows with the aggregate load at the SP. We further summarize the impact of associated SDs on $s^{\mathrm{On}}_{\mathbbm{i}}$ through an effective congestion load defined as $y_{\mathbbm{i}}=\sum_{u^{\mathrm{On}}_{\mathbbm{v}}\in\mathcal{U}^{\mathrm{On}}_{\mathbbm{i}}}\gamma_{\mathbbm{v},\mathbbm{i}}$. With this notation, (10) can be written in a factorized form as:
$t^{\mathrm{comp}}_{\mathbbm{j},\mathbbm{i}}=\theta_{\mathbbm{j},\mathbbm{i}}\;\tau_{\mathbbm{i}}(y_{\mathbbm{i}}),\quad\theta_{\mathbbm{j},\mathbbm{i}}=\frac{r_{\mathbbm{j}}}{\gamma_{\mathbbm{j},\mathbbm{i}}},\quad\tau_{\mathbbm{i}}(y)=\frac{y}{f^{\mathrm{On}}_{\mathbbm{i}}},$ (11)
where $\theta_{\mathbbm{j},\mathbbm{i}}$ is a user-specific load factor, whereas $\tau_{\mathbbm{i}}(y)$ captures the server-side congestion effect as a function of the aggregate load. This separable structure will be instrumental for constructing an ordinal potential in the subsequent analysis.
While (11) characterizes the congestion-induced computing delay, an offloading SD also experiences wireless transmission latency in the uplink: when $u^{\mathrm{On}}_{\mathbbm{j}}$ offloads to $s^{\mathrm{On}}_{\mathbbm{i}}$, the uplink delay is $t^{\mathrm{tx}}_{\mathbbm{j},\mathbbm{i}}=\frac{d_{\mathbbm{j}}}{W\log_{2}\!\big(1+e^{\mathrm{tx}}_{\mathbbm{j}}\,\Gamma_{\mathbbm{j},\mathbbm{i}}\big)}$, where $d_{\mathbbm{j}}$ is the input data size, $W$ is the bandwidth, $e^{\mathrm{tx}}_{\mathbbm{j}}$ is the transmit power, and $\Gamma_{\mathbbm{j},\mathbbm{i}}$ is the Signal-to-Noise-Ratio (SNR) between $u^{\mathrm{On}}_{\mathbbm{j}}$ and $s^{\mathrm{On}}_{\mathbbm{i}}$. Consequently, the total edge-side latency experienced by $u^{\mathrm{On}}_{\mathbbm{j}}$ when uploading its workload to $s^{\mathrm{On}}_{\mathbbm{i}}$ is given by:
$T^{\mathrm{edge}}_{\mathbbm{j},\mathbbm{i}}=t^{\mathrm{comp}}_{\mathbbm{j},\mathbbm{i}}+t^{\mathrm{tx}}_{\mathbbm{j},\mathbbm{i}}.$ (12)
V-A2 Utility of HUs
Building on the above latency and energy models, we now characterize the utility of an HU when deciding whether to execute a task locally or offload it to an SP. Consider an HU $u^{\mathrm{On}}_{\mathbbm{j}}\in\bm{\mathcal{U}}^{\mathrm{H}}$ with local CPU frequency $f^{\mathrm{loc}}_{\mathbbm{j}}$ and local computational power $e^{\mathrm{loc}}_{\mathbbm{j}}$. The local execution time and energy consumption of its task are given by $t^{\mathrm{loc}}_{\mathbbm{j}}={r_{\mathbbm{j}}}/{f^{\mathrm{loc}}_{\mathbbm{j}}}$ and $E^{\mathrm{loc}}_{\mathbbm{j}}={r_{\mathbbm{j}}\,e^{\mathrm{loc}}_{\mathbbm{j}}}/{f^{\mathrm{loc}}_{\mathbbm{j}}}$, respectively. When the HU $u^{\mathrm{On}}_{\mathbbm{j}}$ offloads its task to $s^{\mathrm{On}}_{\mathbbm{i}}$, it experiences latency $t^{\mathrm{comp}}_{\mathbbm{j},\mathbbm{i}}+t^{\mathrm{tx}}_{\mathbbm{j},\mathbbm{i}}$ and uplink energy $E^{\mathrm{tx}}_{\mathbbm{j},\mathbbm{i}}$. Compared with local task execution, this leads to the latency saving of $t^{\mathrm{save}}_{\mathbbm{j},\mathbbm{i}}{=}t^{\mathrm{loc}}_{\mathbbm{j}}-\big(t^{\mathrm{comp}}_{\mathbbm{j},\mathbbm{i}}+t^{\mathrm{tx}}_{\mathbbm{j},\mathbbm{i}}\big)$ and energy saving of $c^{\mathrm{save}}_{\mathbbm{j},\mathbbm{i}}{=}E^{\mathrm{loc}}_{\mathbbm{j}}-E^{\mathrm{tx}}_{\mathbbm{j},\mathbbm{i}}$ ($E^{\mathrm{tx}}_{\mathbbm{j},\mathbbm{i}}$ denotes the uplink transmission energy).
To map these latency and energy savings into a scalar benefit, we introduce non-negative valuation coefficients $\mathbb{V}^{\mathrm{H}}_{t}$ and $\mathbb{V}^{\mathrm{H}}_{e}$ to capture HU sensitivities to latency and energy, respectively. The HU’s perceived valuation of being served by $s^{\mathrm{On}}_{\mathbbm{i}}$ is then modeled via $v^{\mathrm{H}}_{\mathbbm{j},\mathbbm{i}}=\mathbb{V}^{\mathrm{H}}_{t}\,t^{\mathrm{save}}_{\mathbbm{j},\mathbbm{i}}+\mathbb{V}^{\mathrm{H}}_{e}\,c^{\mathrm{save}}_{\mathbbm{j},\mathbbm{i}}$. Given this valuation, the HU’s net utility must also account for the service payment: if $s^{\mathrm{On}}_{\mathbbm{i}}$ charges a price $p_{\mathbbm{j},\mathbbm{i}}$ for serving $u^{\mathrm{On}}_{\mathbbm{j}}$, its net utility is
$U^{\mathrm{H}}_{\mathbbm{j},\mathbbm{i}}=v^{\mathrm{H}}_{\mathbbm{j},\mathbbm{i}}-p_{\mathbbm{j},\mathbbm{i}},\qquad U^{\mathrm{H}}_{\mathbbm{j},0}=0,$ (13)
where the value of $U^{\mathrm{H}}_{\mathbbm{j},0}$ corresponds to local task execution.
V-A3 Utility of MUs
While HUs mainly care about the joint latency–energy tradeoff, MUs (e.g., robots or sensing devices) are typically driven by task deadlines and reliability requirements. We therefore adopt a slightly different utility structure for MUs. Specifically, an MU $u^{\mathrm{On}}_{\mathbbm{j}}\in\bm{\mathcal{U}}^{\mathrm{M}}$ is associated with a latency deadline $\tau^{\mathrm{ddl}}_{\mathbbm{j}}$, while its total edge-side latency at $s^{\mathrm{On}}_{\mathbbm{i}}$ is $T^{\mathrm{edge}}_{\mathbbm{j},\mathbbm{i}}$ in (12). To capture the fact that the MU primarily values meeting this deadline, we adopt a soft-deadline valuation
$G^{\mathrm{M}}_{\mathbbm{j},\mathbbm{i}}=v_{\mathbbm{j}}\,{\big[\tau^{\mathrm{ddl}}_{\mathbbm{j}}-T^{\mathrm{edge}}_{\mathbbm{j},\mathbbm{i}}\big]_{+}}\times({\tau^{\mathrm{ddl}}_{\mathbbm{j}}})^{-1},$ (14)
where $x]_{+}=\max\{x,0\}$ and $v_{\mathbbm{j}}>0$ denotes the MU’s task reward. Intuitively, the MU obtains the full reward $v_{\mathbbm{j}}$ if the deadline is met, whereas excessive latency leads to a loss of reward (once the deadline is violated, the reward becomes zero). In particular, “soft-deadline” means that this reward decreases smoothly as latency approaches the deadline, instead of dropping abruptly from full reward to zero at a hard threshold.
In addition to timeliness, an MU also benefits from saving local computing energy by task offloading. Similar to HUs, the MU’s energy saving when utilizing edge services is $c^{\mathrm{M},\mathrm{save}}_{\mathbbm{j},\mathbbm{i}}=E^{\mathrm{loc}}_{\mathbbm{j}}-E^{\mathrm{tx}}_{\mathbbm{j},\mathbbm{i}}$, where $E^{\mathrm{loc}}_{\mathbbm{j}}$ is the local execution energy and $E^{\mathrm{tx}}_{\mathbbm{j},\mathbbm{i}}$ is the uplink transmission energy. Let $\mathbb{V}^{\mathrm{M}}_{e}$ denote the MU’s valuation coefficient on energy saving. Combining deadline satisfaction, energy saving, and the service payment $p_{\mathbbm{j},\mathbbm{i}}$, the MU’s utility when offloading its tasks to $s^{\mathrm{On}}_{\mathbbm{i}}$ is given by:
$U^{\mathrm{M}}_{\mathbbm{j},\mathbbm{i}}=G^{\mathrm{M}}_{\mathbbm{j},\mathbbm{i}}+\mathbb{V}^{\mathrm{M}}_{e}\,c^{\mathrm{M},\mathrm{save}}_{\mathbbm{j},\mathbbm{i}}-p_{\mathbbm{j},\mathbbm{i}},\qquad U^{\mathrm{M}}_{\mathbbm{j},0}=0,$ (15)
where $U^{\mathrm{M}}_{\mathbbm{j},0}$ corresponds to local task execution.
Note that both HU and MU utilities in (13) and (15) inherently share the same congestion component $t^{\mathrm{comp}}_{\mathbbm{j},\mathbbm{i}}=\theta_{\mathbbm{j},\mathbbm{i}}\tau_{\mathbbm{i}}(y_{\mathbbm{i}})$ in (11). This shared structure will be crucial for aligning individual incentives with system-wide performance via an ordinal potential function in our subsequent analysis (Sec. V-B).
V-A4 Cost of SPs
Having specified how HUs and MUs (i.e., SDs) evaluate edge services, we now focus on the SPs’ perspective. As more SDs offload their tasks to an SP, the SP consumes more energy and experiences higher operating costs, which must be reflected in its payoff and, ultimately, in the system-level SW. Subsequently, for each SP $s^{\mathrm{On}}_{\mathbbm{i}}$, given its aggregated computing load $\mathbbm{x}_{\mathbbm{i}}=\sum_{u^{\mathrm{On}}_{\mathbbm{v}}\in\mathcal{U}^{\mathrm{On}}_{\mathbbm{i}}}r_{\mathbbm{v}}$, we model its computation/operation cost as follows:
$K_{\mathbbm{i}}(\mathbbm{x}_{\mathbbm{i}})=\omega_{1}\,e^{\mathrm{comp}}_{\mathbbm{i}}{\mathbbm{x}_{\mathbbm{i}}}\big/{f^{\mathrm{On}}_{\mathbbm{i}}}+\omega_{2}c^{\mathrm{hard}},$ (16)
where $e^{\mathrm{comp}}_{\mathbbm{i}}$ is the average computing power consumption of $s^{\mathrm{On}}_{\mathbbm{i}}$, and $c^{\mathrm{hard}}$ captures hardware- and maintenance-related costs ($\omega_{1},\omega_{2}\geq 0$ are weighting coefficients). In (16), the term proportional to $\mathbbm{x}_{\mathbbm{i}}/f^{\mathrm{On}}_{\mathbbm{i}}$ reflects that processing a larger workload leads to higher energy and operational expenditure. It is worth noting that although an SP may collect service payments $\sum_{u^{\mathrm{On}}_{\mathbbm{v}}\in\mathcal{U}^{\mathrm{On}}_{\mathbbm{i}}}p_{\mathbbm{v},\mathbbm{i}}$, these payments are monetary transfers among SP and SDs and therefore cancel out when aggregating the system-level SW. Consequently, they do not appear in the potential function constructed later in Sec. V-B and only affect how the SW is divided among SPs and SDs.
In a nutshell, in our considered human–machine ecosystem, SPs resources are limited and SDs (i.e., HUs/MUs) act selfishly, where each SD tries to choose the execution mode of its tasks (local vs. offloading) and SP that maximizes its own utility. Directly coordinating all the coupled decisions across SPs and SDs via a centralized SW maximization formulation is computationally prohibitive and poorly scalable. This motivates our following game-theoretic treatment in which we capture the system performance by a carefully crafted potential function that is aligned with individual incentives and allows for decentralized handling of resource allocations across SPs and SDs.
V-B Ordinal Potential Function
In order to analyze our online resource scheduling game and to enable simple distributed best-response dynamics with provable convergence, we first construct an ordinal potential $\Phi(\bm{\pi})$ that aggregates SD-side benefits and SP-side costs. This construction leverages the unified congestion structure in (11) together with the above-described deadline-aware MUs’ utilities.
For an MU $u^{\mathrm{On}}_{\mathbbm{j}}$ served by $s^{\mathrm{On}}_{\mathbbm{i}}$, the time-related part of its payoff was modeled by the soft-deadline valuation in (14) as $G^{\mathrm{M}}_{\mathbbm{j},\mathbbm{i}}=v_{\mathbbm{j}}\,{\big[\tau^{\mathrm{ddl}}_{\mathbbm{j}}-T^{\mathrm{edge}}_{\mathbbm{j},\mathbbm{i}}\big]_{+}}\times{(\tau^{\mathrm{ddl}}_{\mathbbm{j}})^{-1}}$. This valuation, together with energy saving and the payment term, determines the MU’s overall utility in the online game defined in (15). We note that since the computing delay $t^{\mathrm{comp}}_{\mathbbm{j},\mathbbm{i}}$ in (11), which is inherently encoded in (15), is an increasing function of the congestion load $y_{\mathbbm{i}}$, heavier congestion directly reduces $G^{\mathrm{M}}_{\mathbbm{j},\mathbbm{i}}$ and the utility in (11). This monotone relationship between congestion and MU valuation allows us to align individual best responses with the system-wide potential in the subsequent analysis.
For HUs, the utility in (13) can be decomposed into a congestion-independent valuation part and a congestion-dependent penalty, as follows:
[…]
where $S_{\mathbbm{j}}(\mathbbm{i})$ collects all valuation terms that do not depend on the congestion load $y_{\mathbbm{i}}$, while the second term captures how queueing-induced delay degrades the HU’s utility. Note that the payment $p_{\mathbbm{j},\mathbbm{i}}$ is also independent of $y_{\mathbbm{i}}$, but is written separately because it represents a monetary transfer, which will cancel out at the system level in our subsequent potential analysis.
MU utilities also admit the same structural form: by linearizing $G^{\mathrm{M}}_{\mathbbm{j},\mathbbm{i}}$ around typical delay levels (see Appx. E-D), the MU utility can be rewritten as:
$U^{\mathrm{M}}_{\mathbbm{j},\mathbbm{i}}=\bar{S}_{\mathbbm{j}}(\mathbbm{i})\;-\;\kappa_{\mathbbm{j}}\,\theta_{\mathbbm{j},\mathbbm{i}}\,\tau_{\mathbbm{i}}(y_{\mathbbm{i}})\;-\;p_{\mathbbm{j},\mathbbm{i}},$ (18)
where $\bar{S}_{\mathbbm{j}}(\mathbbm{i})$ gathers the non-congestion valuation terms (soft-deadline valuation base reward and energy saving), and $\kappa_{\mathbbm{j}}>0$ reflects the MU’s effective sensitivity to congestion. Similar to (17), $p_{\mathbbm{j},\mathbbm{i}}$ is separated from $\bar{S}_{\mathbbm{j}}(\mathbbm{i})$ in (18) even though it does not depend on $y_{\mathbbm{i}}$, so that monetary transfers can be excluded from the system-level potential.
Considering both HUs and MUs, the utility of SD $u^{\mathrm{On}}_{\mathbbm{j}}\in\mathcal{U}^{\mathrm{On}}$ when it is served by $s^{\mathrm{On}}_{\mathbbm{i}}\in\{0\}\cup\mathcal{S}^{\mathrm{On}}$ can be unified as follows:
$U_{\mathbbm{j},\mathbbm{i}}=S_{\mathbbm{j}}(\mathbbm{i})\;-\;\kappa_{\mathbbm{j}}\,\theta_{\mathbbm{j},\mathbbm{i}}\,\tau_{\mathbbm{i}}(y_{\mathbbm{i}})\;-\;p_{\mathbbm{j},\mathbbm{i}},$ (19)
where $\mathbbm{i}=0$ corresponds to local task execution at SD $u^{\mathrm{On}}_{\mathbbm{j}}$ and $S_{\mathbbm{j}}(0)=0$. In (19), $S_{\mathbbm{j}}(\mathbbm{i})$ encodes the non-congestion valuation terms (latency/energy saving or soft-deadline valuation), $\kappa_{\mathbbm{j}}\,\theta_{\mathbbm{j},\mathbbm{i}}$ captures how sensitive $u^{\mathrm{On}}_{\mathbbm{j}}$ is to the congestion at $s^{\mathrm{On}}_{\mathbbm{i}}$, and $\tau_{\mathbbm{i}}(\cdot)$ is the server-side congestion function that only applies when $\mathbbm{i}\in\mathcal{S}^{\mathrm{On}}$ (local execution does not impact edge-side congestion). The term $p_{\mathbbm{j},\mathbbm{i}}$ represents the payment charged to $u^{\mathrm{On}}_{\mathbbm{j}}$ and, while independent of congestion, is written explicitly as it will later be treated as a monetary transfer in the system-level analysis. The unified representation in (19) captures SD (i.e., HU/MU) heterogeneity through $(S_{\mathbbm{j}},\kappa_{\mathbbm{j}},\theta_{\mathbbm{j},\mathbbm{i}})$, whereas all SDs are coupled through the same congestion mechanism $\tau_{\mathbbm{i}}(y_{\mathbbm{i}})$ at each SP.
To describe how individual decisions interact at the system-level, we collect all SD assignments into the joint assignment profile [^4] Given $\bm{\pi}$, the binary indicators $x^{\mathrm{On}}_{\mathbbm{j},\mathbbm{i}}$ are induced by $x^{\mathrm{On}}_{\mathbbm{j},\mathbbm{i}}=\mathbbm{1}_{\{\pi_{\mathbbm{j}}=\mathbbm{i}\}},\quad\mathbbm{i}\in\{0\}\cup\mathcal{S}^{\mathrm{On}}$ ($\mathbbm{1}_{\{\cdot\}}$ denotes the binary indicator function), so that the vector $\bm{X}^{\mathrm{On}}_{\mathbbm{j}}$ and the scalar action $\pi_{\mathbbm{j}}$ are equivalent representations of $u^{\mathrm{On}}_{\mathbbm{j}}$’s decision. $\bm{\pi}=(\pi_{\mathbbm{j}})_{\mathbbm{j}\in\mathcal{U}^{\mathrm{On}}}$, where $\pi_{\mathbbm{j}}\in\{0\}\cup\mathcal{S}^{\mathrm{On}}$ denotes the choice of $u^{\mathrm{On}}_{\mathbbm{j}}$: $\pi_{\mathbbm{j}}=0$ means that the task is processed locally at $u^{\mathrm{On}}_{\mathbbm{j}}$, while $\pi_{\mathbbm{j}}=\mathbbm{i}$ indicates task offloading to SP $s^{\mathrm{On}}_{\mathbbm{i}}$. Given a profile $\bm{\pi}$, we can then compute, for each SP $s^{\mathrm{On}}_{\mathbbm{i}}$, the effective congestion load and the physical computing load as follows:
$y_{\mathbbm{i}}(\bm{\pi})=\sum_{\mathbbm{j}:\,\pi_{\mathbbm{j}}=\mathbbm{i}}\gamma_{\mathbbm{j},\mathbbm{i}},\qquad\mathbbm{x}_{\mathbbm{i}}(\bm{\pi})=\sum_{\mathbbm{j}:\,\pi_{\mathbbm{j}}=\mathbbm{i}}r_{\mathbbm{j}},$ (20)
where $y_{\mathbbm{i}}(\bm{\pi})$ feeds into the congestion-induced delay via $t^{\mathrm{comp}}_{\mathbbm{j},\mathbbm{i}}=\theta_{\mathbbm{j},\mathbbm{i}}\,\tau_{\mathbbm{i}}\big(y_{\mathbbm{i}}(\bm{\pi})\big)$ in (11), and $\mathbbm{x}_{\mathbbm{i}}(\bm{\pi})$ determines the SP-side computation/operation cost through $K_{\mathbbm{i}}\big(\mathbbm{x}_{\mathbbm{i}}(\bm{\pi})\big)$ in (16). Thus, any unilateral change of $\pi_{\mathbbm{j}}$ only modifies the loads of the involved SPs in (20), and these updated loads propagate to SD delays and SP costs through (11) and (16), respectively.
In order to analyze the resulting game and to show that simple myopic best-response updates can converge, we now construct an ordinal potential function that aggregates both SD-side benefits and SP-side costs as follows:
$\hskip-5.69054pt{\footnotesize\begin{aligned} \Phi(\bm{\pi}){=}\hskip-19.91692pt&\underbrace{\sum_{\mathbbm{j}:\,\pi_{\mathbbm{j}}\neq 0}S_{\mathbbm{j}}(\pi_{\mathbbm{j}})}_{\text{user-side non-congestion benefits}}\hskip-8.53581pt-\underbrace{\sum_{\mathbbm{i}\in\mathcal{S}^{\mathrm{On}}}\int_{0}^{\,y_{\mathbbm{i}}(\bm{\pi})}\hskip-4.2679pt\tau_{\mathbbm{i}}(z)\,dz}_{\text{aggregated congestion cost}}-\hskip-8.53581pt\underbrace{\sum_{\mathbbm{i}\in\mathcal{S}^{\mathrm{On}}}K_{\mathbbm{i}}\!\big(\mathbbm{x}_{\mathbbm{i}}(\bm{\pi})\big)}_{\text{SP computation/operation cost}}\hskip-5.69054pt,\end{aligned}}\hskip-14.22636pt\vskip 2.84526pt$ (21)
where the first term sums up the non-congestion valuation of all SDs that choose to offload, the second term integrates the congestion function $\tau_{\mathbbm{i}}(\cdot)$ from zero to the realized load $y_{\mathbbm{i}}(\bm{\pi})$ across SPs, and the last term captures the cost of running edge resources. By jointly accounting for SD-side benefits, congestion penalties, and SP-side costs, $\Phi(\bm{\pi})$ serves as a system-level measure whose variations mirror individual utility changes; this coupling is key to proving that our online resource scheduling game is an ordinal potential game and that distributed better/best-response dynamics converge to a pure-strategy NE. In particular, consider a unilateral deviation where one SD $u^{\mathrm{On}}_{\mathbbm{j}}$ switches from SP $s^{\mathrm{On}}_{\mathbbm{i}}$ to another SP $s^{\mathrm{On}}_{\mathbbm{i}^{\prime}}$ (or to local task execution). The resulting change in $\Phi(\bm{\pi})$ can be decomposed as follows:
  • User-side non-congestion term: The first term in (21) only depends on the non-congestion valuation $S_{\mathbbm{j}}(\cdot)$ of the deviating SD, while all other SDs keep the same assignment. Hence, its increment is exactly the change in that SD’s non-congestion benefit, i.e., $S_{\mathbbm{j}}(\pi_{\mathbbm{j}}^{\text{new}})-S_{\mathbbm{j}}(\pi_{\mathbbm{j}}^{\text{old}})$.
  • Congestion term: The second term aggregates the integrated congestion at each SP through $\int_{0}^{y_{\mathbbm{i}}(\bm{\pi})}\tau_{\mathbbm{i}}(z)\,dz$, where $\tau_{\mathbbm{i}}(y)$ is an increasing function of the congestion load $y$ (recall $\tau_{\mathbbm{i}}(y)=y/f^{\mathrm{On}}_{\mathbbm{i}}$). When $u^{\mathrm{On}}_{\mathbbm{j}}$ moves from $s^{\mathrm{On}}_{\mathbbm{i}}$ to $s^{\mathrm{On}}_{\mathbbm{i}^{\prime}}$ (or from/to local execution), the loads $y_{\mathbbm{i}}$ and $y_{\mathbbm{i}^{\prime}}$ change by fixed amounts. Because $\tau_{\mathbbm{i}}(\cdot)$ and $\tau_{\mathbbm{i}^{\prime}}(\cdot)$ are increasing functions and $\kappa_{\mathbbm{j}}\theta_{\mathbbm{j},\mathbbm{i}}>0$, the sign of the change in the integrated congestion term matches the sign of the change in that SD’s congestion penalty $\kappa_{\mathbbm{j}}\theta_{\mathbbm{j},\mathbbm{i}}\tau_{\mathbbm{i}}(y_{\mathbbm{i}})$; this follows the standard integral–pointwise comparison used in congestion games[27].
  • SP-side cost term: The third term in (21) captures the SP-side computation/operation cost $K_{\mathbbm{i}}\big(\mathbbm{x}_{\mathbbm{i}}(\bm{\pi})\big)$, where $\mathbbm{x}_{\mathbbm{i}}(\bm{\pi})$ is the aggregate computing load at $s^{\mathrm{On}}_{\mathbbm{i}}$. A deviation by $u^{\mathrm{On}}_{\mathbbm{j}}$ only changes the loads at $s^{\mathrm{On}}_{\mathbbm{i}}$ and $s^{\mathrm{On}}_{\mathbbm{i}^{\prime}}$, and, due to the convex and increasing behavior of $K_{\mathbbm{i}}(\cdot)$ with respect to $\mathbbm{x}_{\mathbbm{i}}$, the sign of the change in this term is aligned with the way that the SD’s action changes the corresponding SP-side cost.
Putting the above components together, it can be immediately construed that whenever $u^{\mathrm{On}}_{\mathbbm{j}}$ changes its action in a way that strictly increases its own utility (excluding the payment term $p_{\mathbbm{j},\mathbbm{i}}$), the resulting change in $\Phi(\bm{\pi})$ has the same sign as its utility improvement. In other words, $\Phi(\bm{\pi})$ is an ordinal potential for our online resource scheduling game under the soft-deadline valuations: any unilateral utility-improving deviation by an SD strictly increases $\Phi(\bm{\pi})$. This property is crucial as it implies that simple distributed better-/best-response dynamics monotonically climb the potential and therefore converge in finitely many steps to a pure-strategy NE. In particular, we mathematically prove that $\Phi(\bm{\pi})$ in (21) is an ordinal potential, so that any unilateral utility-improving move strictly increases $\Phi(\bm{\pi})$ and any better-/best-response dynamics converge to a pure-strategy NE (see Appx. C for details); the complete technical development, including the matrix reformulation, the integral–pointwise “sandwich” inequality, and the finite-improvement property, is deferred to Appx E.
V-C Problem Formulation
[…]
$\mathrm{s.t.}\ \$$\pi_{\mathbbm{j}}\in\{0\}\cup\mathcal{S}^{\mathrm{On}},\quad\forall\,u^{\mathrm{On}}_{\mathbbm{j}}\in\mathcal{U}^{\mathrm{On}},$ (22a)
$x_{\mathbbm{j},\mathbbm{i}}^{\mathrm{On}}=0,\ \ \text{if }s^{\mathrm{On}}_{\mathbbm{i}}\notin\bm{C}_{\mathbbm{j}},\quad\forall\,u^{\mathrm{On}}_{\mathbbm{j}}\in\mathcal{U}^{\mathrm{On}},$ (22b)
$\sum_{u^{\mathrm{On}}_{\mathbbm{j}}\in{\mathcal{U}^{\mathrm{On}}}}x_{\mathbbm{j},\mathbbm{i}}^{\mathrm{On}}\leq G_{\mathbbm{i}},\quad\forall\,s^{\mathrm{On}}_{\mathbbm{i}}\in\mathcal{S}^{\mathrm{On}},$ (22c)
where $\bm{\pi}=(\pi_{\mathbbm{j}})_{\mathbbm{j}\in\mathcal{U}^{\mathrm{On}}}$ denotes the joint assignment profile, with $\pi_{\mathbbm{j}}=0$ representing local task execution at $u^{\mathrm{On}}_{\mathbbm{j}}$ and $\pi_{\mathbbm{j}}=\mathbbm{i}$ indicating task offloading to SP $s^{\mathrm{On}}_{\mathbbm{i}}$. The binary variables $x_{\mathbbm{j},\mathbbm{i}}^{\mathrm{On}}$ are induced by $\bm{\pi}$ as defined in Sec. V-B (i.e., $x_{\mathbbm{j},\mathbbm{i}}^{\mathrm{On}}=1$ if $\pi_{\mathbbm{j}}=\mathbbm{i}$).
Constraint (22a) enforces each SD to choose only one execution option for its current task: either local processing ($\pi_{\mathbbm{j}}=0$) or offloading to a single SP ($\pi_{\mathbbm{j}}\in\mathcal{S}^{\mathrm{On}}$). Constraint (22b) captures the service-coverage feasibility: $\bm{C}_{\mathbbm{j}}$ denotes the set of SPs that can physically serve $u^{\mathrm{On}}_{\mathbbm{j}}$ (e.g., within communication range), and thus the association variable for an SD outside the coverage of $s^{\mathrm{On}}_{\mathbbm{i}}$ is forced to zero. Constraint (22c) limits the admission load at each SP, where $G_{\mathbbm{i}}$ is the maximum number of SDs that SP $s^{\mathrm{On}}_{\mathbbm{i}}$ can concurrently serve.
[…]
V-D Potential-Guided Best-Response Dynamics (PG-BRD)
For brevity, we provide the complete pseudo-code of PG-BRD in Appx. D and summarize its four main steps below:
Step 1. Random-feasible initialization: We first construct a feasible initial association profile $\bm{\pi}^{(0)}\in\{0,1,\ldots,|\mathcal{S}^{\mathrm{On}}|\}^{|\mathcal{U}^{\mathrm{On}}|}$, where $\pi_{\mathbbm{j}}=0$ corresponds to local task execution at $u^{\mathrm{On}}_{\mathbbm{j}}$ and $\pi_{\mathbbm{j}}=\mathbbm{i}>0$ means that $u^{\mathrm{On}}_{\mathbbm{j}}$ offloads its task to SP $s^{\mathrm{On}}_{\mathbbm{i}}$. For each SD $u^{\mathrm{On}}_{\mathbbm{j}}$, we randomly select an action from the set $\{0\}\cup\bm{C}_{\mathbbm{j}}$ and only accept it if the resulting profile satisfies the coverage constraint (22b) and the subcarrier-capacity constraint (22c). This procedure ensures that the algorithm starts from a valid resource scheduling configuration with no constraint violations.
Step 2. Feasible-action enumeration under coverage and subcarrier-capacity constraints: Given a current profile $\bm{\pi}$, we compute for each SP $s^{\mathrm{On}}_{\mathbbm{i}}$ the effective congestion load $y_{\mathbbm{i}}(\bm{\pi})$ and the physical computing load $\mathbbm{x}_{\mathbbm{i}}(\bm{\pi})$ as in (20), as well as the number of currently served SDs $n_{\mathbbm{i}}(\bm{\pi})=\big|\{\,\mathbbm{j}\mid\pi_{\mathbbm{j}}=\mathbbm{i}\,\}\big|$. When it is SD $u^{\mathrm{On}}_{\mathbbm{j}}$’s turn to update its decision, we temporarily remove its own contribution and obtain the loads generated by all other SDs, denoted by $y^{-{\mathbbm{j}}}_{\mathbbm{i}}$, $\mathbbm{x}^{-{\mathbbm{j}}}_{\mathbbm{i}}$, and $n^{-{\mathbbm{j}}}_{\mathbbm{i}}$. These quantities represent the congestion level and computing load at each SP if $u^{\mathrm{On}}_{\mathbbm{j}}$ were not associated with any SP. Given these congestion and load levels generated by all other SDs, we construct the feasible action set for $u^{\mathrm{On}}_{\mathbbm{j}}$ as follows:
$\mathcal{A}_{\mathbbm{j}}^{\mathrm{feas}}{=}\Big\{a\in\{0\}\cup\mathcal{S}^{\mathrm{On}}\hskip 0.0pt\Big|\hskip 0.0pta=0\ \text{or}\ \big(a=\mathbbm{i}>0,\ s^{\mathrm{On}}_{\mathbbm{i}}\in\bm{C}_{\mathbbm{j}},\ n^{-{\mathbbm{j}}}_{\mathbbm{i}}+1\leq G_{\mathbbm{i}}\big)\Big\},$
where $a=0$ corresponds to local task execution, and an offloading action $a=\mathbbm{i}>0$ is feasible only if SP $s^{\mathrm{On}}_{\mathbbm{i}}$ covers $u^{\mathrm{On}}_{\mathbbm{j}}$ and still has enough residual subcarrier-capacity to admit one more SD without violating its capacity $G_{\mathbbm{i}}$. This step explicitly removes any candidate action that would break either the coverage or the subcarrier-capacity constraints.
Step 3. Asynchronous best response under the ordinal potential: In each iteration, SDs are visited one-by-one in a random order, and at most one SD is allowed to change its decision at a time. For a given SD $u^{\mathrm{On}}_{\mathbbm{j}}$ and current profile $\bm{\pi}$, we first evaluate its current utility $U_{\mathbbm{j}}(\pi_{\mathbbm{j}}\mid\bm{\pi}_{-\mathbbm{j}})$ using the unified expression in (19). Then, for each feasible action $a\in\mathcal{A}_{\mathbbm{j}}^{\mathrm{feas}}$, we temporarily set $\pi_{\mathbbm{j}}=a$, recompute the congestion and delay quantities for this SD, and obtain the resulting utility $U_{\mathbbm{j}}(a\mid\bm{\pi}_{-\mathbbm{j}})$. The SD then chooses a best response $a^{\mathrm{best}}_{\mathbbm{j}}\in\arg\max_{a\in\mathcal{A}_{\mathbbm{j}}^{\mathrm{feas}}}U_{\mathbbm{j}}(a\mid\bm{\pi}_{-\mathbbm{j}})$. To avoid frequent switches caused by marginal gains, we introduce an acceptance threshold $\varepsilon>0$ and let $u^{\mathrm{On}}_{\mathbbm{j}}$ update its decision only if $U_{\mathbbm{j}}(a^{\mathrm{best}}_{\mathbbm{j}}\mid\bm{\pi}_{-\mathbbm{j}})-U_{\mathbbm{j}}(\pi_{\mathbbm{j}}\mid\bm{\pi}_{-\mathbbm{j}})>\varepsilon$. Otherwise, it keeps its current action $\pi_{\mathbbm{j}}$ unchanged. Because the game admits the ordinal potential $\Phi(\bm{\pi})$, every accepted update in this step strictly increases $\Phi(\bm{\pi})$.
Step 4. Convergence and resulting equilibrium: The above two steps (feasible-action enumeration and asynchronous best responses) are repeatedly applied by going over all SDs. When a complete pass over $\mathcal{U}^{\mathrm{On}}$ produces no accepted update, no SD can improve its utility by more than $\varepsilon$ through a unilateral change of its action, and thus a pure-strategy NE $\bm{\pi}^{\star}$ is reached.
VI Evaluations
[…]
VI-A Benchmark Methods and Evaluation Metrics
We evaluate FUSION from two complementary angles: (i) prediction performance on service demand forecasting module (Sec. VI-B); and (ii) system-level performance (Sec. VI-C).
VI-A1 Benchmarks and Metrics for Demand Forecasting
We compare Pro-LNN against two widely used prediction models.
$\bullet$ LSTM-driven demand prediction[28]: A standard long short-term memory (LSTM) network is applied to the input sequences $\{\mathbf{z}_{i}\}$ and trained to predict future ES demands. This baseline represents a strong discrete-time RNN model.
$\bullet$ Transformer-driven demand prediction[29]: A temporal Transformer-based predictor with self-attention is used over the historical context window. This serves as a powerful baseline, but typically requires more computing resources.
Performance evaluations consider the following metrics:
$\bullet$ Accuracy: We report the root mean square error (RMSE), mean absolute error (MAE), and symmetric mean absolute percentage error (sMAPE) of the $H$-step forecasts $\{\hat{N}_{i}^{\mathrm{Dem},(\nu+1:\nu+H)}\}$ against the ground truth $\{N_{i}^{\mathrm{Dem},(\nu+1:\nu+H)}\}$, averaged over all ESs and all test instances.
$\bullet$ Model footprint: We measure the total number of trainable parameters and the corresponding 32-bit floating-point format (FP32, measured in megabytes) for each predictor, which together characterize its computing (memory/storage) overhead.
$\bullet$ Robustness: To assess robustness under realistic noisy conditions, we compare the mean squared error (MSE) on clean test inputs with that on corrupted inputs containing randomly missing observations, additive perturbations, and occasional demand bursts, and report the performance drop $\Delta\text{MSE}$.
VI-A2 System-Level Benchmarks and Metrics
We then compare FUSION with several benchmark methods enumerated below:
$\bullet$ Pure online scheduling (PurOnline)[13]: This baseline removes the offline stage of FUSION completely, e.g., ESs and APs do not sign any offline contracts, and UAV locations are not pre-planned. All HUs and MUs directly enter the online market, where their tasks are scheduled only via the PG-BRD-based scheduling over the current ES/AP network environment.
$\bullet$ Two-stage resource provisioning without PG-based online scheduling (FUSION_NoPG): This variant keeps the same offline stage as FUSION, i.e., Off-AIC2 with eACO-VRP is still used to form offline AP–ES contracts and pre-planned routes. However, the online stage abandons PG-BRD: HU/MU tasks are randomly assigned to feasible ESs/APs within coverage subject to resource capacity and deadline constraints.
$\bullet$ Two-stage resource provisioning with random offline contracts (FUSION_Random): This baseline uses the same online stage as FUSION and still relies on PG-BRD for congestion-aware online resource scheduling. However, its offline stage no longer uses Off-AIC2 or eACO-VRP; instead, each AP is randomly associated with a subset of ESs and moves along a random feasible route until its capacity is exhausted.
$\bullet$ Two-stage resource provisioning without ES spatio-temporal features (FUSION_NoST): This baseline preserves the two-stage structure of FUSION, but its offline auction ignores ES-specific time windows and travel-time constraints. Specifically, AP-ES pairs are determined only by simple price/volume trading without using eACO-VRP, while online stage still runs PG-BRD for task scheduling.
TABLE II: Robustness–complexity comparison of different forecasters on synthetic non-stationary series. The “corrupt” case injects 30% random missing values, Gaussian noise with standard deviation $\sigma=0.05$, and random burst shocks.
Model #Params $\text{MSE}_{\text{clean}}$ $\text{MSE}_{\text{corrupt}}$ $\Delta\text{MSE}$ Rel. increase
LSTM 52800 0.119 0.983 0.864 +726.3%
Transformer 69500 0.467 0.781 0.314 +67.3%
Pro-LNN 11100 0.459 0.762 0.303 +66.1%
To assess performance, we use the following metrics:
$\bullet$ Social welfare (SW): SW captures the sum of utilities of all ESs, APs, and users across both stages: in the offline stage, SW accounts for ES profits, AP utilities, and auctioneer surplus given by (1)-(3); in the online stage, it captures the potential value $\Phi(\bm{\pi})$, combining SD-side benefits and SP-side costs.
$\bullet$ Delay incurred by interactions between SPs and SDs (DoI): DoI captures the total decision-making latency (in milliseconds) induced by control-message exchanges between ESs/APs (SPs) and HUs/MUs (SDs) across both offline and online stages. To compute DoI, following [30, 32, 33], we assume that the per-interaction delay lies in the range of $[1,15]$ ms.
$\bullet$ Energy consumption incurred by interactions between SPs and SDs (ECoI): ECoI measures the signaling-related energy consumption during the decision-making process. To compute ECoI, we assume that the transmit power of ESs lies in $[6,20]$ W, consistent with [30, 32]. Specifically, ECoI is obtained by aggregating the energy consumed by all control-message transmissions and the corresponding interaction delays.
$\bullet$ Truthfulness and individual rationality of Off-AIC2: To empirically verify the economic properties of Off-AIC2, we conduct a set of experiments in which ESs and APs are allowed to deviate from their truthful bids/asks. We measure the proportion of ESs and APs whose utilities increase under such deviations. An auction that is truthful and individually rational should ensure that (i) every winning ES attains a non-negative utility and cannot benefit from misreporting its bid, and (ii) every winning AP receives a payment no lower than its declared ask and gains no advantage by inflating or deflating it.
VI-B Performance Evaluation on Prediction
VI-B1 Robustness vs. Cost
We first conduct experiments to evaluate “robustness vs. cost” trade-off in Table II. In the “clean” setting, we feed each model with intact historical sequences, while in the “corrupt” setting, we inject 30% random missing values, additive Gaussian noise with standard deviation $\sigma=0.05$, and random burst shocks of length 1–3. On clean data, LSTM attains the lowest MSE (0.119), whereas Transformer and Pro-LNN yield slightly larger errors (0.467 and 0.459, respectively). This gap is expected, as LSTM uses approximately 52.8k parameters, compared to only 11.1k parameters of Pro-LNN, implying significantly higher computation costs. Once corrupted inputs are introduced, however, LSTM error surges to 0.983 (a relative increase of +726.3%), while the errors of Transformer and Pro-LNN only grow to 0.781 and 0.762 (relative increases of +67.3% and +66.1%), respectively. These results indicate that our proposed Pro-LNN offers a favorable trade-off between robustness and cost compared to an over-parameterized LSTM on noisy sequences, which matches the characteristics of ES demand trajectories in Off-AIC2.
VI-B2 Real-World Dataset Test
[…]
TABLE III: Prediction accuracy and model footprint on the UCI dataset (feeders MT_005 and MT_145, input length $T_{\mathrm{his}}=48$, prediction horizon $H=10$).
Model #Params Size (MB) RMSE MAE sMAPE
LSTM 52842 0.204 0.1121 0.0816 0.1831
Transformer 69482 0.400 0.1158 0.0824 0.1864
Pro-LNN 11114 0.046 0.1234 0.0906 0.1983
TABLE IV: Robustness to missing and noisy inputs on UCI Electricity (30% random missing values, added Gaussian noise with standard deviation of $\sigma=0.05$, and random burst shocks).
[…]
LSTM 0.0126 0.0485 0.0359
Transformer 0.0134 0.0592 0.0458
Pro-LNN 0.0152 0.0467 0.0315
Figure 2: Performance comparisons and economic property analyses, where (a)-(b) SW under different problem sizes (80 HUs, 80 MUs, and 5 APs in (a), and 400 HUs, 400 MUs, and 35 APs in (b)), (c)-(d) DoI and ECoI under different problem sizes (s1-s6 are set as $\{75/75/15/8\}$, $\{75/75/20/8\}$, $\{90/90/20/10\}$, $\{90/90/25/10\}$, $\{115/115/25/12\}$, and $\{115/115/30/12\}$), (e)-(f) truthfulness, and (g)-(h) individual rationality.
VI-C Performance Evaluation on FUSION
To emulate heterogeneous, time-varying online demands across different ESs, we adopt the real-world UCI Electricity Load Diagrams dataset [34], select several distinct load profiles, and map their normalized power-consumption traces to the per-time-slot resource demand of each ES. We then train Pro-LNN on these profiles to forecast the per-ES demand [^5] We employ the UCI dataset as a trace to capture realistic temporal patterns of aggregate demand at each ES. Pro-LNN is trained on these time series and used to forecast the demand $N_{i}^{\mathrm{Dem},(\nu)}$ of ES $s_{i}$ over future trading rounds; the forecasted demand profiles drive the time-varying SD task-arrival intensities at each ES, while the total population sizes of SDs and SPs are kept fixed to control the overall network scale., and then evaluate FUSION as follows (i.e., in Sec. VI-C1–Sec. VI-C3).
VI-C1 Social Welfare (SW)
Figs. 22 report SW as the number of ESs increases from 8 to 16. The two subfigures correspond to two representative network configurations: 80 HUs, 80 MUs, and 5 APs in Fig. 2, and 400 HUs, 400 MUs, and 35 APs in Fig. 2 to capture different network scales.
In Fig. 2, SW increases for all methods as more ESs are introduced, since additional ESs provide more computing resources and service opportunities. However, the growth rates differ across schemes. FUSION consistently achieves the highest SW because its offline stage employs Off-AIC2 to match AP resources with ESs of high revenue potential under spatio-temporal constraints, while its online stage adopts a PG-BRD scheduler to perform congestion-aware task allocation among HUs and MUs, leading to more efficient and fair resource utilization. PurOnline, which lacks an offline stage and relies solely on an online potential game over real-time ES resources, achieves the lowest SW due to the absence of proactive UAV repositioning. FUSION_NoPG shares the same offline design as FUSION but replaces PG-BRD with random online assignment, resulting in unresolved congestion and a notable SW degradation, highlighting the importance of the online PG. FUSION_Random adopts random AP–ES association in the offline stage while retaining PG-BRD online, yielding intermediate performance and indicating that intelligent online scheduling alone cannot compensate for the lack of revenue- and spatio-temporal-aware offline planning. FUSION_NoST ignores ES spatio-temporal characteristics in the offline stage and relies on a price-driven auction with PG-BRD online; it outperforms PurOnline and FUSION_Random but still exhibits a consistent gap to FUSION. Fig. 2 shows qualitatively similar trends under the larger-scale configuration: as the number of ESs increases, SW rises for all methods, with FUSION consistently outperforming the baselines. Overall, these results indicate that, in FUSION, a well-designed offline AP–ES contracting mechanism (particularly, Off-AIC2 with explicit spatio-temporal and routing considerations) effectively unlocks scalability gains, while the PG-BRD-based online stage mitigates congestion and further improves resource utilization.
VI-C2 Evaluation on Interaction Overhead
To quantitatively assess interaction efficiency, we analyze DoI (Fig. 2) and ECoI (Fig. 2). The x-axis (s1–s6) denotes different system configurations (various HU/MU/ES/AP combinations detailed in the caption of the figure), and the y-axis uses a logarithmic scale to capture results across magnitudes. Across all settings, DoI and ECoI remain within the same order of magnitude: FUSION is not always the best in DoI/ECoI, but the gaps with baselines (FUSION_NoPG, FUSION_Random, FUSION_NoST, PurOnline) are rather small. This arises from FUSION’s offline contracting stage, which exposes more candidate SPs and resource options, prompting the PG-BRD online scheduler to explore a larger strategy space, slightly increasing interaction rounds and per-transaction overhead. Importantly, FUSION’s auction and route planning occur offline, adding no extra delay or energy cost to the online stage (DoI/ECoI in Figs. 22 reflect amortized offline overhead). Combined with SW results (Figs. 2-2), we conclude that FUSION maintains comparable DoI/ECoI to baselines while substantially improving SW through more efficient air-ground resource utilization.
VI-C3 Truthfulness and Individual Rationality
To assess truthfulness and individual rationality, we study the Off-AIC2 embedded in FUSION, under a representative setting with 75 HUs, 75 MUs, 12 ESs, and 7 APs. Figs. 2 and 2 show the utilities of sellers (APs) and buyers (ESs) as functions of their reported asks/bids, while Figs. 2 and 2 verify individual rationality by comparing true costs/values with payments.
Fig. 2 shows four independent market instances, where one AP is randomly selected in each subplot. The horizontal axis denotes the reported ask and the vertical axis the resulting utility (revenue minus true cost), with the blue dot marking the true ask. As the ask increases, the AP’s utility rises while it remains selected, but drops sharply once the ask exceeds its critical region and the contract is lost. In all instances, the utility is maximized near the true ask, indicating that cost misreporting does not improve the seller’s payoff. Fig. 2 presents a similar analysis for ESs. In each subplot, one ES is selected, with the horizontal axis representing the reported mean bid and the red dot indicating the true valuation. When the bid is below a threshold close to the true value, the ES fails to win and obtains zero utility; once this threshold is crossed, the ES becomes a winner and gains utility, while further bid increases yield negligible gains. Across instances, the best-response region consistently concentrates around the true valuation. The above near-truthful behavior is due to the critical-value-based pricing adopted in the third stage of Off-AIC2. A binary-search procedure is used to approximate each participant’s critical price, i.e., the lowest effective bid for an ES or the highest effective ask for an AP that guarantees winning. Payments are then determined according to the classical critical-value pricing principle, under which large deviations from true valuations tend to exclude participants from the winner set. Since the allocation problem jointly couples ES–AP matching and route planning and is solved heuristically via eACO-VRP, strict dominant-strategy truthfulness cannot be guaranteed; nevertheless, the empirical results demonstrate consistent near-truthful behavior.
The negative utilities observed in parts of Figs. 2 and 2 correspond to extreme hypothetical misreports. They arise because utilities are evaluated using true valuations while reported bids/asks are artificially searched over a wide range; such loss-making strategies would not be chosen by rational participants in practice. Figs. 2 and 2 further confirm individual rationality. For randomly selected winning APs, the total reward determined by Off-AIC2 is always no smaller than the corresponding true total cost, and for selected ESs, the true valuation always exceeds the final payment.
VII Conclusion and Future Work
In this work, we proposed FUSION, a forecasting-driven two-stage framework for service provisioning in air–ground integrated networks with human–machine coexistence. In its offline stage, FUSION leverages Pro-LNN to capture multi-step spatio-temporal dynamics of ES workloads and integrates the predictions into eACO-VRP and Off-AIC2 to jointly plan AP routes and design incentive-compatible ES–AP contracts. In its online stage, congestion-aware task scheduling is formulated as an ordinal potential game over heterogeneous HUs, MUs, and SPs (ESs and APs), and addressed via PG-BRD algorithm that obtains a pure-strategy NE while satisfying coverage and capacity constraints. Experiments on both synthetic and real-world datasets show that FUSION consistently achieves higher social welfare and improved resource utilization, while maintaining latency and energy costs comparable to state-of-the-art baselines and preserving individual rationality, budget balance, and near-truthfulness.
Promising future directions include integrating quantum-powered ESs and investigating resource-efficient training and inference of large-scale foundation models in UAV-assisted air–ground networks.
Appendix A Key Notations
Key notations in this paper are summarized in Table V.
TABLE V: Key notations
Notation Explanation
$\bm{\mathcal{S}}$, $\bm{\mathcal{U}}^{\mathrm{H}}$, $\bm{\mathcal{U}}^{\mathrm{M}}$, $\bm{\mathcal{V}}$ Sets of edge servers (ESs), human users (HUs), machine users (MUs), and air–ground providers (APs)
$s_{i}$, $u_{m}^{\mathrm{H}}$, $u_{n}^{\mathrm{M}}$, $v_{k}$ The $i$-th ES, the $m$-th HU, the $n$-th MU, and the $k$-th AP
$N_{i}^{\mathrm{Dem},(\nu)}$, $\hat{N}_{i}^{\mathrm{Dem},(\nu+h)}$ Actual and LNN-predicted task demand of ES $s_{i}$ at trading rounds $\nu$ and $\nu\!+\!h$
$l_{i}^{\mathrm{E}}$, $l_{m}^{\mathrm{H}}$, $l_{n}^{\mathrm{M}}$, $l_{k}^{\mathrm{A}}$ 3D locations of ES $s_{i}$, HU $u_{m}^{\mathrm{H}}$, MU $u_{n}^{\mathrm{M}}$, and the dispatching vehicle of AP $v_{k}$
$r_{m}^{\mathrm{H}}$, $r_{n}^{\mathrm{M}}$, $r_{\mathbbm{j}}$ Computation workload (CPU cycles) of HU $u_{m}^{\mathrm{H}}$, MU $u_{n}^{\mathrm{M}}$, and generic SD $u^{\mathrm{On}}_{\mathbbm{j}}$
$d_{m}^{\mathrm{H}}$, $d_{n}^{\mathrm{M}}$, $d_{\mathbbm{j}}$ Input data size (bits) of HU tasks, MU tasks, and generic SD tasks
$f_{i}^{\mathrm{E}}$, $f_{\mathbbm{i}}^{\mathrm{On}}$, $f_{m}^{\mathrm{H,loc}}$, $f_{n}^{\mathrm{M,loc}}$, $f_{k}^{\mathrm{A}}$ Computing capabilities of ES $s_{i}$, online SP $s_{\mathbbm{i}}^{\mathrm{On}}$, local HU/MU devices, and UAVs dispatched by AP $v_{k}$
$e_{m}^{\mathrm{H,tx}}$, $e_{n}^{\mathrm{M,tx}}$ Uplink transmit power of HU $u_{m}^{\mathrm{H}}$ and MU $u_{n}^{\mathrm{M}}$
$e_{i}^{\mathrm{E}}$ Local computing energy consumption of ES $s_{i}$
$e_{k}^{\mathrm{comp}}$, $e_{k}^{\mathrm{move}}$, $e_{k}^{\mathrm{fly}}$ Energy consumption of AP $v_{k}$ for computation, ground movement, and UAV flight
$\tau_{n}^{\mathrm{ddl}}$, $\tau^{\mathrm{ddl}}_{\mathbbm{j}}$ Task deadline of MU $u_{n}^{\mathrm{M}}$ in offline stage and latency deadline of SD $u^{\mathrm{On}}_{\mathbbm{j}}$ in online stage
$C_{k}^{\mathrm{A}}$ Total number of UAVs carried by AP $v_{k}$
$N_{i,k}^{\mathrm{Trad}}$ Quantity of traded computing resources between ES $s_{i}$ and AP $v_{k}$ in an offline contract
$x_{i,k}^{\mathrm{Off}}$, $\bm{X}$ Binary offline assignment between ES $s_{i}$ and AP $v_{k}$; $\bm{X}$ collects all $x_{i,k}^{\mathrm{Off}}$
$p_{i,k}^{\mathrm{ES}}$, $r_{i,k}^{\mathrm{ES}}$, $\widetilde{p}_{i}$ Payment from ES $s_{i}$ to AP $v_{k}$, reward received by $v_{k}$, and historical average user service price at $s_{i}$
$U_{i}^{\mathrm{ES,Off}}$, $U_{k}^{\mathrm{A}}$, $U^{\mathrm{Auc}}$ Utility of ES $s_{i}$, utility of AP $v_{k}$, and utility of the auctioneer in offline stage
$U^{\mathrm{SW}}$ Total SW of the offline auction market (sum of utilities of ESs, APs, and auctioneer)
[…]
$\varphi_{r,s}$, $\eta_{r,s}$, $p^{\kappa}_{\mathrm{n}_{r}\rightarrow\mathrm{n}_{s}}$ Pheromone level, distance, and transition probability on edge $(\mathrm{n}_{r},\mathrm{n}_{s})$ in eACO-VRP
$\mathcal{S}^{\mathrm{On}}$, $\mathcal{U}^{\mathrm{On}}$, $\mathcal{H}$, $\mathcal{A}$ Sets of online service providers (SPs), service demanders (SDs), HUs and MUs in online stage
$x_{\mathbbm{j}}^{\mathrm{On}}$, $x_{\mathbbm{j},\mathbbm{i}}^{\mathrm{On}}$, $\mathbf{x}_{\mathbbm{j}}$ Indicator of local execution, indicator of scheduling SD $u^{\mathrm{On}}_{\mathbbm{j}}$ to SP $s^{\mathrm{On}}_{\mathbbm{i}}$, and the scheduling vector of SD $u^{\mathrm{On}}_{\mathbbm{j}}$
$y_{\mathbbm{i}}$, $\mathbbm{x}_{\mathbbm{i}}$, $n_{\mathbbm{i}}(\bm{\pi})$ Effective congestion load, aggregate computing load, and number of associated SDs at SP $s^{\mathrm{On}}_{\mathbbm{i}}$ under profile $\bm{\pi}$
$t^{\mathrm{comp}}_{\mathbbm{j},\mathbbm{i}}$, $t^{\mathrm{tx}}_{\mathbbm{j},\mathbbm{i}}$, $T^{\mathrm{edge}}_{\mathbbm{j},\mathbbm{i}}$ Computation delay, uplink transmission delay, and total edge-side latency of SD $u^{\mathrm{On}}_{\mathbbm{j}}$ at SP $s^{\mathrm{On}}_{\mathbbm{i}}$
$S_{\mathbbm{j}}(\mathbbm{i})$, $\kappa_{\mathbbm{j}}$, $\theta_{\mathbbm{j},\mathbbm{i}}$, $\tau_{\mathbbm{i}}(y)$ Non-congestion benefit term, congestion sensitivity, load factor, and congestion function at SP $s^{\mathrm{On}}_{\mathbbm{i}}$
$K_{\mathbbm{i}}(\mathbbm{x}_{\mathbbm{i}})$, $C_{\mathbbm{i}}$ Computation/operation cost function and maximum computing capacity of SP $s^{\mathrm{On}}_{\mathbbm{i}}$
$\bm{\pi}$, $\Phi(\bm{\pi})$, $\bm{\mathcal{F}}^{(\mathrm{I})}, 𝓕$$\bm{\mathcal{F}}^{(\mathrm{II})}$ Joint scheduling profile, ordinal potential function, and optimization problems in offline stage and online stage
Appendix B Detailed of Offline Stage
B-A Interaction Flow for Forecasting, Deployment, and Contracting
The overall interaction flow of the proposed Pro-LNN + eACO-VRP + Off-AIC2 framework as follow:
$\bullet$ At the beginning of each offline trading round, each ES locally runs Pro-LNN to predict its service demand over future rounds, and submits the resulting demand profile as bid information to the auction platform. Each AP, in turn, reports its available service capacity and ask price based on its computing capability, energy consumption, and available service duration.
$\bullet$ After collecting all bids and asks, the auction platform first invokes Off-AIC2 to filter out infeasible ES–AP pairs under price and capacity constraints, and then calls eACO-VRP [^6] Note that eACO-VRP serves as a route-aware matching subroutine inside Off-AIC2, and the two methods are tightly coupled in the actual mechanism. For clarity of exposition, however, we present the standalone design of eACO-VRP before introducing the complete Off-AIC2 procedure later in this subsection. on the resulting candidate set to determine, for each winning AP, the set of ESs to be served and their visiting order, thereby generating pre-planned routes that jointly account for time windows, travel distance, and utility.
$\bullet$ Subsequently, Off-AIC2 applies its critical-price-based pricing rules to compute settlement payments for winning ESs and rewards for winning APs.
$\bullet$ The final offline outcome includes the ES–AP matching relations, the payment vectors, and the pre-planned service routes for each AP, which will be taken as inputs to the online stage over the subsequent time horizon.
B-B Detailed Pro-LNN Architecture and Training Procedure
In this section, we provide the detailed update equations and training procedure of the proposed Pro-LNN predictor used in Off-AIC2.
Step 1. Stable liquid cell (line 7, Alg. 1): At the core of Pro-LNN is a stable liquid cell that updates a hidden state $\mathbf{h}_{i}^{(\nu)}\in\mathbb{R}^{d}$ given the current input $\mathbf{z}_{i}^{(\nu)}$ and the previous state. Concretely, we first compute neuron-wise time constants and a pre-activation: $\bm{\tau}^{(\nu)}=\mathrm{softplus}\big(W_{\tau}[\mathbf{z}_{i}^{(\nu)},\mathbf{h}_{i}^{(\nu)}]\big)+\tau_{\min}$ and $\tilde{\mathbf{u}}^{(\nu)}=W_{h}\mathbf{h}_{i}^{(\nu)}+W_{x}\mathbf{z}_{i}^{(\nu)}+\mathbf{b}$, where $W_{\tau},W_{h},W_{x}$ and $\mathbf{b}$ are trainable, and $\tau_{\min}>0$ enforces strictly positive time scales (we also clamp $\bm{\tau}^{(\nu)}$ to $\tau_{\min},\tau_{\max}]$ for numerical stability). The effective update rate is then $\bm{\alpha}^{(\nu)}=\mathrm{clip}\Big(\frac{\Delta\nu}{\bm{\tau}^{(\nu)}},\alpha_{\min},\alpha_{\max}\Big)$, with a fixed $\Delta\nu$ (one trading interval in our implementation) and small constants $0<\alpha_{\min}\leq\alpha_{\max}<1$. We then apply a lightweight layer normalization and non-linearity, $\tilde{\mathbf{h}}_{i}^{(\nu)}=\tanh\!\big(\mathrm{LN}(\tilde{\mathbf{u}}^{(\nu)})\big)$, and update the hidden state as
$\mathbf{h}_{i}^{(\nu+1)}=(1-\bm{\alpha}^{(\nu)})\odot\mathbf{h}_{i}^{(\nu)}+\bm{\alpha}^{(\nu)}\odot\tilde{\mathbf{h}}_{i}^{(\nu)}+\gamma\,\tanh\!\big(\mathrm{LN}(\mathbf{h}_{i}^{(\nu)})\big),$ (24)
where $\gamma$ is a small residual scaling factor and $\odot$ denotes element-wise multiplication. This leaky-integrator update blends exponential memory decay with input-driven state injection, while the neuron-wise $\bm{\tau}^{(\nu)}$ allow different neurons to operate at different effective time scales.
Notation. For clarity, all nonlinear operators in the above update are applied element-wise unless otherwise stated. Specifically, $\mathrm{softplus}(\cdot)$ denotes the standard softplus activation $\mathrm{softplus}(x)=\log(1+e^{x})$, $\mathrm{clip}(x,a,b)$ clips each element of $x$ into the interval $[a,b]$, $\mathrm{LN}(\cdot)$ denotes layer normalization [24], and $\mathrm{LiquidCell}(\mathbf{z}_{i}^{(\nu)},\mathbf{h}_{i}^{(\nu)})$ is a shorthand for one step of the stable liquid update in Step 1 (using the time-constant gating and residual integration described above).
Step 2. Historical encoding (lines 4-7, Alg. 1): Given the latest $L$ observations of ES $s_{i}$ up to trading round $\nu$, Pro-LNN encodes the history by iteratively applying the liquid cell: $\mathbf{h}_{i}^{(\nu-L+1)}\leftarrow\mathrm{LiquidCell}\big(\mathbf{z}_{i}^{(\nu-L+1)},\mathbf{0}\big),\ldots,\mathbf{h}_{i}^{(\nu)}\leftarrow\mathrm{LiquidCell}\big(\mathbf{z}_{i}^{(\nu)},\mathbf{h}_{i}^{(\nu-1)}\big)$.
This exponential-style integration naturally adapts memory decay and information injection under irregular or noisy trajectories.
Step 3. Multi-horizon readout and consistency (lines 8-17, Alg. 1): After encoding the history, a shallow MLP readout maps the final hidden state to an $H$-step forecast:
$\hskip-8.53581pt\hat{\mathbf{n}}_{i}^{(1:H)}=\mathrm{Readout}\big(\mathbf{h}_{i}^{(\nu)}\big)=[\hat{N}_{i}^{\mathrm{Dem},(\nu+1)},\ldots,\hat{N}_{i}^{\mathrm{Dem},(\nu+H)}],$ (25)
where $\mathrm{Readout}(\cdot)$ denotes a lightweight fully connected network that maps the $d$-dimensional hidden state to an $H$-dimensional prediction vector. The base training loss is the mean squared error (MSE) over the $H$ horizons. To further improve stability for long horizons, we optionally add a multi-step consistency objective: starting from a history window, we repeatedly roll the model forward in closed loop by feeding its own one-step predictions back into the window, and penalize the mismatch at each rolled-out step.
Input: Training sequences $\{\mathbf{z}_{i}^{(1:T)}\}$; history length $L$; horizon $H$; learning rate $\eta$; consistency depth $K$
Output: Predicted sequences $\{\hat{N}_{i}^{\mathrm{Dem},(\nu+1:\nu+H)}\}$
1
2Initialize liquid-cell and readout parameters $\theta$;
3 while not converged do
4  Sample a mini-batch $\mathcal{B}$ of training windows from $\mathcal{D}$;
5  for each window $(\mathbf{z}_{i}^{(\nu-L+1:\nu)},N_{i}^{\mathrm{Dem},(\nu+1:\nu+H)})\in\mathcal{B}$ do
6     Set hidden state $\mathbf{h}_{i}\leftarrow\mathbf{0}$;
7     for $k=\nu-L+1$ to $\nu$ do
8        $\mathbf{h}_{i}\leftarrow\mathrm{LiquidCell}\big(\mathbf{z}_{i}^{(k)},\mathbf{h}_{i}\big)$;
9    (Base $H$-step forecast)
10     $\hat{\mathbf{n}}_{i}^{(1:H)}\leftarrow\mathrm{Readout}(\mathbf{h}_{i})$;
11     $\mathcal{L}_{i}\leftarrow\frac{1}{H}\sum_{h=1}^{H}\big(\hat{n}_{i}^{(h)}-N_{i}^{\mathrm{Dem},(\nu+h)}\big)^{2}$;
12     if $K>0$ then
13        Construct a history window $\mathbf{y}_{i}$ from $\{\mathbf{z}_{i}^{(\nu-L+1)},\ldots,\mathbf{z}_{i}^{(\nu)}\}$;
14        for $k=1$ to $K$ do
15           $\hat{\mathbf{n}}_{i}^{(1:H)}\leftarrow f_{\theta}(\mathbf{y}_{i})$;
16           $\mathcal{L}_{i}\leftarrow\mathcal{L}_{i}+\big(\hat{n}_{i}^{(1)}-N_{i}^{\mathrm{Dem},(\nu+k)}\big)^{2}$;
17           Update $\mathbf{y}_{i}$ by discarding the oldest step and appending $\hat{n}_{i}^{(1)}$ as the newest demand;
18       $\mathcal{L}_{i}\leftarrow\mathcal{L}_{i}/(H+K)$;
19
20 Aggregate loss $\mathcal{L}\leftarrow\frac{1}{|\mathcal{B}|}\sum_{i\in\mathcal{B}}\mathcal{L}_{i}+\lambda\,\Omega(\theta)$;
21  Update parameters $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}$;
Algorithm 1 Pro-LNN
Step 4. Inference and integration (line 9, Alg. 1): At inference, Pro-LNN takes, for each ES $s_{i}$, the latest $L$ observations $\{\mathbf{z}_{i}^{(\nu-L+1)},\ldots,\mathbf{z}_{i}^{(\nu)}\}$, encodes them via the liquid cell into $\mathbf{h}_{i}^{(\nu)}$, and applies the readout to obtain an $H$-step forecast $\{\hat{N}_{i}^{\mathrm{Dem},(\nu+1)},\ldots,\hat{N}_{i}^{\mathrm{Dem},(\nu+H)}\}$. These predictions are computed in the offline stage before each trading round and then fed into: (i) the eACO-VRP module for time-window-aware ES recommendation and AP route pre-planning, and (ii) the auction module for ES–AP contract design. Since Pro-LNN runs offline, its wall-clock inference latency does not affect the real-time trading between SDs and SPs, while its compact footprint and robustness to noisy inputs make it well suited for deployment on edge and near-edge computing platforms.
B-C Detailed pseudo-code of eACO-VRP
In this section, we provide the detailed pseudo-code of eACO-VRP in Alg. 2.
[…]
2
[…]
4
5while $iter\leq iter_{\mathrm{max}}$ do
6  for each ant $\mathbbm{k}\in\mathbb{K}$ do
7     $\bm{M}^{\mathbbm{k}}\!\leftarrow\!\varnothing$;  $\mathrm{n}_{r}\!\leftarrow\!\mathrm{n}_{0}$
8     while end condition not met do
9        Compute: feasible set $J^{\mathbbm{k}}(\mathrm{n}_{r})$ and transition probabilities $pr^{\mathbbm{k}}_{r,s}$ (via (6))
10        if $J^{\mathbbm{k}}(\mathrm{n}_{r})\neq\varnothing$ then
11           sample next node $\mathrm{n}_{s}$ according to $p^{\mathbbm{k}}_{r,s}$ from $J^{\mathbbm{k}}(\mathrm{n}_{r})$
12           $\bm{M}^{\mathbbm{k}}\!\leftarrow\!\bm{M}^{\mathbbm{k}}\cup\{\text{edge}(\mathrm{n}_{r},\mathrm{n}_{s})\}$
13           $\mathrm{n}_{r}\!\leftarrow\!\mathrm{n}_{s}$
14       else
15          break
16
17
18
19 $\bm{M}^{\mathrm{max}}\!\leftarrow\!\arg\max_{\bm{M}^{\mathbbm{k}}:\mathbbm{k}\in\mathbb{K}}\;\sum_{(\mathrm{n}_{r},\mathrm{n}_{s})\in\bm{M}^{\mathbbm{k}}}\triangle{U^{\mathrm{W}}_{r,s}}$
[…]
21    Pheromone evaporation and deposit (7) and (8)
22
23 $iter\!\leftarrow\!iter+1$
24
Output: $\bm{M}^{\mathrm{max}}$ (pre-planned route for $v_{k}$), and the recommended ES tasks along $\bm{M}^{\mathrm{max}}$
Algorithm 2 Proposed eACO-VRP
B-D Detailed pseudo-code of Off-AIC2
In this section, we provide the detailed pseudo-code of Off-AIC2 in Alg. 3.
1Input: ES set $\bm{\mathcal{S}}$ with demands $\{N_{i}^{\mathrm{Dem}}\}$ and per-ES bids $\{\overline{\mathrm{bid}}_{i}\}$; AP set $\bm{\mathcal{V}}$ with asks $\{\mathrm{ask}_{k}\}$ and capacities $\{C_{k}^{\mathrm{A}}\}$
2Initialization: Set $\bm{X}^{*}\!\leftarrow\!\bm{0}$, $\bm{p}^{\mathrm{b}*}\!\leftarrow\!\bm{0}$, $\bm{r}^{\mathrm{s}*}\!\leftarrow\!\bm{0}$.
3Sort ES indices into $\mathcal{L}^{\mathrm{ES}}$ by $\overline{\mathrm{bid}}_{i}$ (descending), and AP indices into $\mathcal{L}^{\mathrm{AP}}$ by $\mathrm{ask}_{k}$ (ascending); relabel $(i,k)$ accordingly.
4 Find a pair of pivotal ranks $(I^{\ast},K^{\ast})$ such that $\overline{\mathrm{bid}}_{I^{\ast}}\geq\mathrm{ask}_{K^{\ast}},\quad\sum_{k=1}^{K^{\ast}}C_{k}^{\mathrm{A}}\;\geq\;\sum_{i=1}^{I^{\ast}}N_{i}^{\mathrm{Dem}}$, and $I^{\ast}$ is maximized (breaking ties by larger capacity slack $\sum_{k=1}^{K^{\ast}}C_{k}^{\mathrm{A}}-\sum_{i=1}^{I^{\ast}}N_{i}^{\mathrm{Dem}}$).
5Initialize the set of unmatched ESs: $\mathcal{U}^{\mathrm{free}}\leftarrow\{s_{i}\mid 1\leq i\leq I^{\ast}\}$.
6 for $k$ in top-$K^{\ast}$ APs (according to $\mathcal{L}^{\mathrm{AP}}$) do
7  if $C_{k}^{\mathrm{A}}\leq 0$ or $\mathcal{U}^{\mathrm{free}}=\emptyset$ then
8     continue
9 Build candidate ES set $\mathcal{C}_{k}\leftarrow\big\{s_{i}\in\mathcal{U}^{\mathrm{free}}\,\big|\,U_{i,k}^{\mathrm{Off}}>0\ \wedge\ \overline{\mathrm{bid}}_{i}\geq\mathrm{ask}_{k}\big\}$;
10 if $\mathcal{C}_{k}=\emptyset$ then
11     continue
12 $\mathcal{K}_{k}^{\star}\leftarrow$ Alg. 2 ;
13  for $i\in\mathcal{K}_{k}^{\star}$ do
14     Set $x^{\mathrm{Off}}_{i,k}\leftarrow 1$;
15     Remove $s_{i}$ from $\mathcal{U}^{\mathrm{free}}$: $\mathcal{U}^{\mathrm{free}}\leftarrow\mathcal{U}^{\mathrm{free}}\setminus\{s_{i}\}$;
16 Update remaining capacity: $C_{k}^{\mathrm{A}}\leftarrow C_{k}^{\mathrm{A}}-\sum_{i\in\mathcal{K}_{k}^{\star}}N_{i}^{\mathrm{Dem}}$.
17Denote the resulting matching by $\bm{X}^{*}=\{x^{\mathrm{Off}}_{i,k}\}$.
18for each matched ES $s_{i}$ with $\sum_{k}x^{\mathrm{Off}}_{i,k}=1$ do
19  Set $high\leftarrow\overline{\mathrm{bid}}_{i}$, $low\leftarrow\overline{\mathrm{bid}}_{I^{\ast}+1}$;
20  while $low<high$ do
21     Set $b_{\mathrm{eff}}\leftarrow(low+high)/2$;
22     Rerun Step 2 on the top-$I^{\ast}$ ESs and top-$K^{\ast}$ APs, using $b_{\mathrm{eff}}$ as ES $s_{i}$’s bid and keeping all other bids/asks unchanged, to obtain a temporary matching $\bm{X}^{\mathrm{Temp}}$;
23     if $\sum_{k}x^{\mathrm{Temp}}_{i,k}=1$ then $high\leftarrow b_{\mathrm{eff}}$;
24     else $low\leftarrow b_{\mathrm{eff}}$;
25
26 Set $p_{i}^{\mathrm{b}*}\leftarrow high$.
27
28for each matched AP $v_{k}$ with $\sum_{s_{i}}x^{\mathrm{Off}}_{i,k}\geq 1$ do
29  Set $high\leftarrow\mathrm{ask}_{k}$, $low\leftarrow\mathrm{ask}_{K^{\ast}+1}$ (or a safe reserve);
30  while $low<high$ do
31     Set $\mathrm{ask}_{\mathrm{eff}}\leftarrow(low+high)/2$;
32     Rerun Step 2 on the top-$I^{\ast}$ ESs and top-$K^{\ast}$ APs, using $\mathrm{ask}_{\mathrm{eff}}$ as AP $k$’s ask (and keeping all other bids/asks unchanged), to obtain a temporary matching $\bm{X}^{\mathrm{Temp}}$;
33     if $\sum_{i}x^{\mathrm{Temp}}_{i,k}\geq 1$ then $low\leftarrow\mathrm{ask}_{\mathrm{eff}}$;
34     else $high\leftarrow\mathrm{ask}_{\mathrm{eff}}$;
35
36 Set $r_{k}^{\mathrm{s}*}\leftarrow low$.
37
Output: Matching $\bm{X}^{*}$, ES payments $\bm{p}^{\mathrm{b}*}$, AP rewards $\bm{r}^{\mathrm{s}*}$.
Algorithm 3 Proposed Off-AIC2
B-E Economic Properties of Off-AIC2
Here, we discuss the fundamental properties that underpin the design of auction in offline stage[14, 22].
Proposition 1.
(Individual Rationality) An auction is said to be individually rational if the payment made by any winning ES (to its matched AP) does not exceed its bid, and the revenue obtained by any winning AP is no less than its asked price.
Proposition 2.
(Near-Truthfulness) An auction mechanism is said to be near-truthful if every ES and AP maximizes, or almost maximizes, its utility by reporting its true private valuation. In particular, for any fixed bids and asks of the other participants, no single ES or AP can increase its utility by more than a small margin through unilaterally misreporting its bid or ask. In this sense, truthful reporting remains an empirically near-optimal strategy, and large deviations from true valuations are not beneficial.
Proposition 3.
(Budget Balance) An auction satisfies budget balance if the total payment collected from all winning ESs is no less than the total payment made to all winning APs.
Proposition 4.
All winning ESs and APs obtained from Off-AIC2 are individually rational during the offline stage.
The proposed offline stage of FUSION is shown to meet all the above properties, with detailed theoretical analysis as follow.
Proof.
The individual rationality of Off-AIC2 is guaranteed by the core design of its allocation (Step 2) and pricing (Step 3) procedures in Alg. 3, under the single-parameter assumption that each ES $s_{i}$ has a per-unit valuation and each AP $v_{k}$ has a per-unit cost.
(i) ES side (buyers). We first consider ES buyers $s_{i}$.
$\bullet$ Allocation in Step 2: For each AP $v_{k}$, the candidate ES set is constructed as
$\mathcal{C}_{k}=\big\{i\in\mathcal{U}^{\mathrm{free}}\,\big|\,U_{i,k}>0\ \wedge\ \mathrm{bid}_{i,k}\geq\overline{\mathrm{bid}}_{i}\big\},$ (26)
where $U_{i,k}=\mathrm{bid}_{i,k}-\mathrm{ask}_{k}$ denotes the surplus of matching ES $s_{i}$ with AP $v_{k}$. Hence, if ES $s_{i}$ is eventually matched (i.e., there exists some $k$ such that $x^{\mathrm{Off}}_{i,k}=1$), the corresponding pair $(s_{i},v_{k})$ must satisfy
$U_{i,k}=\mathrm{bid}_{i,k}-\mathrm{ask}_{k}>0,\qquad\mathrm{bid}_{i,k}\geq\overline{\mathrm{bid}}_{i}.$ (27)
Under truthful bidding, $\mathrm{bid}_{i,k}$ equals $s_{i}$’s true per-unit valuation for the contracted service. Therefore, already at the allocation stage, every winning ES enjoys a strictly positive surplus margin between its valuation and the AP’s ask.
$\bullet$ Pricing in Step 3: In Step 3, for each winning ES $s_{i}$ with $\sum_{k}x^{\mathrm{Off}}_{i,k}=1$, Off-AIC2 treats its effective bid as a variable and performs a monotone binary search over an interval of the form
$p_{i}^{\mathrm{b}*}\in[\,\underline{b}_{i},\overline{\mathrm{bid}}_{i}\,],$ (28)
where $\overline{\mathrm{bid}}_{i}$ is $s_{i}$’s reference bid (coinciding with its true valuation when truthful) and $\underline{b}_{i}$ is a market-wide lower bound. The binary search stops at the minimal critical payment that keeps $s_{i}$ in the winning set when Step 2 is re-executed. By construction, we have
$p_{i}^{\mathrm{b}*}\leq\overline{\mathrm{bid}}_{i}.$ (29)
Moreover, for the realized winning pair $(i,k)$ we also have $\mathrm{bid}_{i,k}\geq\overline{\mathrm{bid}}_{i}$. Combining these relations yields
$\mathrm{bid}_{i,k}\;\geq\;\overline{\mathrm{bid}}_{i}\;\geq\;p_{i}^{\mathrm{b}*}.$ (30)
Under truthful bidding, $\mathrm{bid}_{i,k}$ is ES $s_{i}$’s true valuation, while $p_{i}^{\mathrm{b}*}$ is its actual per-unit payment in Off-AIC2. Hence the ex-post utility of $s_{i}$ in the offline stage satisfies
$u_{i}^{\mathrm{ES}}=\mathrm{(valuation)}-\mathrm{(payment)}\;\geq\;\mathrm{bid}_{i,k}-p_{i}^{\mathrm{b}*}\;\geq\;0,$ (31)
which proves individual rationality for all winning ESs.
(ii) AP side (sellers). We now turn to AP sellers $v_{k}$.
$\bullet$ Allocation in Step 2: In Step 2, APs only serve ESs that yield strictly positive surplus, and they are protected by capacity constraints. Specifically, $v_{k}$ may serve ES $s_{i}$ only if $U_{i,k}=\mathrm{bid}_{i,k}-\mathrm{ask}_{k}>0$, and the matching obeys
$\sum_{i}x^{\mathrm{Off}}_{i,k}N_{i}^{\mathrm{Dem}}\leq C_{k}^{\mathrm{A}}.$ (32)
Under truthful reporting, $\mathrm{ask}_{k}$ equals $v_{k}$’s true per-unit cost. Thus each winning AP only accepts contracts where the buyer-side valuation exceeds its cost, ensuring a nonnegative surplus margin at the allocation stage.
$\bullet$ Pricing in Step 3: In Step 3, for each winning AP $v_{k}$ with $\sum_{i}x^{\mathrm{Off}}_{i,k}\geq 1$, Off-AIC2 performs a monotone binary search over an interval
$r_{k}^{\mathrm{s}*}\in[\,\mathrm{ask}_{k},\overline{r}_{k}\,],$ (33)
where $\mathrm{ask}_{k}$ is its own ask (true cost when truthful) and $\overline{r}_{k}$ is a market-determined upper bound. The search stops at the maximal critical reward that still keeps $v_{k}$ in the winning set when Step 2 is rerun. Consequently, we have
$r_{k}^{\mathrm{s}*}\;\geq\;\mathrm{ask}_{k}.$ (34)
Under truthful reporting, $\mathrm{ask}_{k}$ is $v_{k}$’s true per-unit cost, and $r_{k}^{\mathrm{s}*}$ is the actual per-unit reward it receives. Hence the ex-post utility of $v_{k}$ in the offline stage is
$u_{k}^{\mathrm{AP}}=\big(r_{k}^{\mathrm{s}*}-\mathrm{ask}_{k}\big)\cdot\Big(\sum_{i}x^{\mathrm{Off}}_{i,k}N_{i}^{\mathrm{Dem}}\Big)\;\geq\;0.$ (35)
In summary, by construction of the candidate sets and critical-value pricing, every winning ES pays no more than its valuation, and every winning AP receives no less than its cost. Therefore, Off-AIC2 is individually rational for both ESs and APs in the offline stage. ∎
Proposition 5 (Near-truthfulness of Off-AIC2).
Off-AIC2 mechanism is near-truthful for both ESs and APs in the offline stage.
Proof.
At a high level, Off-AIC2 follows the same structural design as classical truthful single-parameter auctions: on both the ES side and the AP side, it combines a (approximately) monotone allocation rule with critical-value-like payments obtained via binary search. In fully monotone settings, such a combination is sufficient for dominant-strategy incentive compatibility; in our case, the presence of route planning via eACO-VRP makes the allocation only approximately monotone, and therefore we obtain approximate (rather than exact) truthfulness.
(i) ES side (buyers). Fix any ES $s_{i}$ and treat all other bids and asks as given. Under the single-parameter simplification, $s_{i}$ reports a single per-unit bid $b_{i}$, which induces its bid row $\mathrm{bid}_{i,k}=b_{i}$ and average bid $\overline{\mathrm{bid}}_{i}=b_{i}$.
Approximate monotone allocation (Steps 1 & 2). In Step 1, ESs are sorted in nonincreasing order of $\overline{\mathrm{bid}}_{i}=b_{i}$. Increasing $b_{i}$ can only move $s_{i}$ upward in this list; it never pushes $s_{i}$ to a worse rank. Since Off-AIC2 only considers the top-$I^{\ast}$ ESs, once $s_{i}$ belongs to the top-$I^{\ast}$ at some bid $b_{i}$, it remains in the top-$I^{\ast}$ for any higher bid $b^{\prime}_{i}>b_{i}$.
In Step 2, a necessary condition for $s_{i}$ to be matched is that it lies in the top-$I^{\ast}$ ESs and that there exists an AP $v_{k}$ such that
$U_{i,k}^{\mathrm{Off}}=\mathrm{bid}_{i,k}-\mathrm{ask}_{k}=b_{i}-\mathrm{ask}_{k}>0,\qquad\overline{\mathrm{bid}}_{i}\geq\mathrm{ask}_{k}.$ (36)
If we ignore the internal randomness and heuristic nature of eACO-VRP, raising $b_{i}$ strictly increases all $U_{i,k}^{\mathrm{Off}}$, making $s_{i}$ (weakly) more attractive and therefore easier to be selected. In practice, the eACO-VRP-based matching can occasionally break strict monotonicity because ES–AP matching and route planning are coupled in a combinatorial way; however, for a fixed tie-breaking rule and fixed routes of other APs, the allocation rule behaves approximately monotone in $b_{i}$: once $s_{i}$ is selected at some bid level, it typically remains selected for any sufficiently close higher bid $b^{\prime}_{i}>b_{i}$.
Critical-value-like payment (Step 3). In Step 3, for each winning ES $s_{i}$ (with $\sum_{k}x^{\mathrm{Off}}_{i,k}=1$), Off-AIC2 treats its effective bid as a variable and performs a monotone binary search over an interval
$b_{i}^{\mathrm{eff}}\in[\,\underline{b}_{i},\overline{\mathrm{bid}}_{i}\,],$ (37)
repeatedly re-running Step 2 to check whether $s_{i}$ would still be selected. The search stops at the smallest bid at which $s_{i}$ remains a winner, and the resulting value is taken as its payment $p_{i}^{\mathrm{b}*}$. When the allocation subroutine is exactly monotone, $p_{i}^{\mathrm{b}*}$ coincides with the classical critical value and yields exact dominant-strategy truthfulness. With eACO-VRP, the binary-search procedure still emulates such a critical value and produces a payment that depends only weakly on $s_{i}$’s own report as long as $b_{i}$ stays in a small neighborhood above $p_{i}^{\mathrm{b}*}$. Our numerical results (Fig. 2) confirm that $s_{i}$’s realized utility curve is single-peaked around its true valuation, and deviations from truthful bidding only bring marginal utility gains, if any. Hence truthful bidding is a locally optimal strategy and nearly utility-maximizing for ESs.
(ii) AP side (sellers). Now fix an AP $v_{k}$ and consider its per-unit ask $a_{k}$, treating all ES bids and other APs’ asks as fixed. Under the single-parameter simplification, $a_{k}$ is $v_{k}$’s single strategic parameter.
Approximate monotone allocation (Steps 1 & 2). In Step 1, APs are sorted in nondecreasing order of $\mathrm{ask}_{k}=a_{k}$. Lowering $a_{k}$ can only move $v_{k}$ upward in this list, never downward. Since Off-AIC2 only considers the top-$K^{\ast}$ APs, once $v_{k}$ lies in the top-$K^{\ast}$ at some ask $a_{k}$, it remains in the top-$K^{\ast}$ for any lower ask $a^{\prime}_{k}<a_{k}$.
In Step 2, $v_{k}$’s attractiveness to ESs is determined by the surplus terms
$U_{i,k}^{\mathrm{Off}}=\mathrm{bid}_{i,k}-a_{k}.$ (38)
Lowering $a_{k}$ strictly increases all $U_{i,k}^{\mathrm{Off}}$, making $v_{k}$ more competitive. With a fixed eACO-VRP configuration, this behavior again yields an allocation rule that is approximately monotone in $a_{k}$: once $v_{k}$ is selected at some ask $a_{k}$, it tends to remain selected for any sufficiently close lower ask $a^{\prime}_{k}<a_{k}$, except for rare non-monotone cases introduced by heuristic route decisions.
Critical-reward-like payment (Step 3). For each winning AP $v_{k}$ (with $\sum_{i}x^{\mathrm{Off}}_{i,k}\geq 1$), Step 3 treats its effective ask as a variable and performs a binary search over
$a_{k}^{\mathrm{eff}}\in[\,\mathrm{ask}_{K^{\ast}+1},\mathrm{ask}_{k}\,],$ (39)
re-running Step 2 to check whether $v_{k}$ would still be selected. The search terminates at the largest ask at which $v_{k}$ remains a winner, and the corresponding reward $r_{k}^{\mathrm{s}*}$ plays the role of a critical ask. When the allocation is exactly monotone, this would give a classical critical payment; with eACO-VRP, $r_{k}^{\mathrm{s}*}$ still behaves as a critical-reward approximation. Empirically (Fig. 2), $v_{k}$’s realized utility as a function of $a_{k}$ exhibits a clear peak near its true ask, and over-reporting beyond that region mostly increases the risk of losing the contract without improving utility. Thus, truthful asking is again a locally best or near-best strategy for APs.
Putting the buyer and seller sides together, Off-AIC2 combines (i) an allocation rule that is monotone up to small perturbations due to the heuristic eACO-VRP subroutine, and (ii) payments and rewards that closely emulate critical values via binary search. According to our definition of approximate truthfulness, this implies that, for both ESs and APs, truthful reporting is a (local) utility-maximizing strategy and any unilateral deviation from the true valuation changes the utility by at most a small margin, as also corroborated by the numerical evidence in Figs. 22. Therefore, Off-AIC2 is near-truthful for both ESs and APs in the offline stage. ∎
Proposition 6.
Off-AIC2 satisfies the (weak) budget-balance property in the offline stage.
Proof.
To verify budget balance, it suffices to show that the auctioneer’s utility in the offline stage,
$U^{\mathrm{P}}=\sum_{s_{i}\in\bm{\mathcal{S}}}\sum_{v_{k}\in\bm{\mathcal{V}}}x^{\mathrm{Off}}_{i,k}N^{\mathrm{{Trad}}}_{i,k}\big(p^{\mathrm{{ES}}}_{i,k}-r_{i,k}^{\mathrm{ES}}\big),$ (40)
is nonnegative, i.e., $U^{\mathrm{P}}\geq 0$. Since $N^{\mathrm{{Trad}}}_{i,k}\geq 0$ and $x^{\mathrm{Off}}_{i,k}\in\{0,1\}$, this reduces to proving that for every active trading pair $(i,k)$ with $x^{\mathrm{Off}}_{i,k}=1$,
$p^{\mathrm{{ES}}}_{i,k}-r_{i,k}^{\mathrm{ES}}\;\geq\;0.$ (41)
(i) Phase 1 (pivotal-rank selection). In Step 1 of Alg. 3, ESs are sorted in nonincreasing order of their average bids $\overline{\mathrm{bid}}_{i}$, and APs are sorted in nondecreasing order of their asks $\mathrm{ask}_{k}$. The pivotal ranks $(I^{\ast},K^{\ast})$ are chosen such that
$\overline{\mathrm{bid}}_{I^{\ast}}\;\geq\;\mathrm{ask}_{K^{\ast}},\qquad\sum_{k=1}^{K^{\ast}}C_{k}^{\mathrm{A}}\;\geq\;\sum_{i=1}^{I^{\ast}}N_{i}^{\mathrm{Dem}}.$ (42)
Alg. 3 only considers the top-$I^{\ast}$ ESs and top-$K^{\ast}$ APs for matching. Hence, for any winning pair $(i,k)$ with $x^{\mathrm{Off}}_{i,k}=1$, we have
$\mathrm{bid}_{i,k}\;\geq\;\overline{\mathrm{bid}}_{I^{\ast}}\;\geq\;\mathrm{ask}_{K^{\ast}}\;\geq\;\mathrm{ask}_{k}.$ (43)
Equation (43) shows that, for each matched pair $(i,k)$, there exists a nonempty price band
$\big\{(p,r)\,\big|\,\mathrm{ask}_{k}\leq r\leq p\leq\mathrm{bid}_{i,k}\big\}$ (44)
in which both ES $s_{i}$ and AP $v_{k}$ obtain nonnegative utilities and the auctioneer’s per-unit margin prp-r is nonnegative.
(ii) Phase 3 (critical pricing). Step 3 of Alg. 3 determines the actual payments within this feasible band using critical-value binary searches:
$\bullet$ Buyer side (ESs): For each winning ES $s_{i}$ with $\sum_{k}x^{\mathrm{Off}}_{i,k}=1$, the payment $p_{i}^{\mathrm{b}*}$ is obtained via a monotone binary search on its effective bid parameter. The search interval is chosen inside the buyer side of the surplus-feasible band in (43), and the algorithm stops at the minimum critical bid that keeps $s_{i}$ in the winning set when Step 2 is re-executed. Thus, for any matched pair $(i,k)$,
$p^{\mathrm{{ES}}}_{i,k}\;=\;p_{i}^{\mathrm{b}*}\;\leq\;\mathrm{bid}_{i,k}.$ (45)
$\bullet$ Seller side (APs): Similarly, for each winning AP $v_{k}$ with $\sum_{i}x^{\mathrm{Off}}_{i,k}\geq 1$, the reward $r_{k}^{\mathrm{s}*}$ is obtained via a monotone binary search on its effective ask parameter. The search interval is chosen inside the seller side of the band in (43), and the algorithm stops at the maximum critical ask that keeps $v_{k}$ in the winning set. Hence, for any matched pair $(i,k)$,
$r_{i,k}^{\mathrm{ES}}\;=\;r_{k}^{\mathrm{s}*}\;\geq\;\mathrm{ask}_{k}.$ (46)
Because both searches are confined to the common surplus-feasible band identified in (43), off-AIC2 always chooses prices satisfying
$\mathrm{ask}_{k}\;\leq\;r_{i,k}^{\mathrm{ES}}\;\leq\;p^{\mathrm{{ES}}}_{i,k}\;\leq\;\mathrm{bid}_{i,k},\qquad\forall(i,k)\ \text{with }x^{\mathrm{Off}}_{i,k}=1.$ (47)
In particular, we obtain
$p^{\mathrm{{ES}}}_{i,k}-r_{i,k}^{\mathrm{ES}}\;\geq\;0\qquad\forall(i,k)\ \text{with }x^{\mathrm{Off}}_{i,k}=1.$ (48)
(iii) Concluding budget balance. Therefore, each active trading pair contributes a nonnegative margin
$x^{\mathrm{Off}}_{i,k}N^{\mathrm{{Trad}}}_{i,k}\big(p^{\mathrm{{ES}}}_{i,k}-r_{i,k}^{\mathrm{ES}}\big)\;\geq\;0,$ (49)
while unmatched pairs contribute zero. Summing over all ES–AP pairs gives
$U^{\mathrm{P}}=\sum_{i,k}x^{\mathrm{Off}}_{i,k}N^{\mathrm{{Trad}}}_{i,k}\big(p^{\mathrm{{ES}}}_{i,k}-r_{i,k}^{\mathrm{ES}}\big)\;\geq\;0,$ (50)
which completes the proof that Off-AIC2 is (weakly) budget balanced in the offline stage. ∎
Appendix C Properties of the Ordinal Potential
To turn the above game formulation into a practical distributed scheduling mechanism, we need a global scalar function whose improvement is aligned with individual SD incentives. In particular, if $\Phi(\bm{\pi})$ acts as an ordinal potential, then any unilateral utility-improving move by an SD strictly increases $\Phi(\bm{\pi})$. This property is exactly what allows simple asynchronous better-/best-response dynamics to converge to a pure-strategy NE with provable guarantees and low online complexity. We therefore establish that the function $\Phi(\bm{\pi})$ in (21) is an ordinal potential for the online game under soft-deadline valuations.
The proof follows a standard single-user deviation argument and relies on three ingredients: (i) separability of the non-congestion terms $S_{\mathbbm{j}}(\cdot)$, (ii) an integral–pointwise comparison that links the pointwise congestion penalty in user utilities with the integrated congestion cost in $\Phi(\bm{\pi})$, and (iii) monotonicity and convexity of the SP-side cost $K_{\mathbbm{i}}(\cdot)$. Full details are deferred to Appendices E-CE-G.
Proposition 7 (Ordinal potential).
Consider the unified utilities in (19) with soft-deadline MUs and the potential $\Phi(\bm{\pi})$ in (21). For any SD $u^{\mathrm{On}}_{\mathbbm{j}}$ and any unilateral deviation $\pi_{\mathbbm{j}}=\mathbbm{i}\to\mathbbm{i}^{\prime}$, if $U_{\mathbbm{j},\mathbbm{i}^{\prime}}-U_{\mathbbm{j},\mathbbm{i}}>0$, then $\Phi\big(\bm{\pi}_{-\mathbbm{j}},\mathbbm{i}^{\prime}\big)-\Phi\big(\bm{\pi}_{-\mathbbm{j}},\mathbbm{i}\big)>0$. Hence, $\Phi(\bm{\pi})$ is an ordinal potential for the considered game[13].
Since each SD possesses a finite action set, any sequence of unilateral best-response (or better-response) updates generates a strictly increasing trajectory of the potential function $\Phi(\bm{\pi})$ and therefore converges in finitely many iterations to a pure-strategy NE. Equivalently, every local maximizer of $\Phi(\bm{\pi})$ corresponds to a pure-strategy NE of the online scheduling game. A formal proof of this finite-improvement property is provided in Appx. E-H.
Appendix D Detailed pseudo-code of PG-BRD
In this section, we provide the detailed pseudo-code of PG-BRD in Alg. 4.
1Input: SD set $\mathcal{U}^{\mathrm{On}}$, SP set $\mathcal{S}^{\mathrm{On}}$, coverage sets $\{\bm{C}_{\mathbbm{j}}\}$, subcarrier-capacity $\{G_{\mathbbm{i}}\}$, max iterations $T_{\max}$, threshold $\varepsilon$.
2[2pt]
3Initialization: Construct a random feasible profile $\bm{\pi}=(\pi_{\mathbbm{j}})_{\mathbbm{j}\in\mathcal{U}^{\mathrm{On}}}$ that satisfies constraints (22a)–(22c); set $\mathbbm{t}\leftarrow 0$ and $\mathrm{converged}\leftarrow\textbf{false}$.
4[2pt]
5while not $\mathrm{converged}$ and $\mathbbm{t}<T_{\max}$ do
6
7 $\mathrm{converged}\leftarrow\textbf{true}$;
8  Generate a random permutation $\mathcal{P}$ of indices in $\mathcal{U}^{\mathrm{On}}$;
9
10 for each SD $u^{\mathrm{On}}_{\mathbbm{j}}$ in $\mathcal{P}$ do
11     let $a^{\mathrm{cur}}_{\mathbbm{j}}\leftarrow\pi_{\mathbbm{j}}$;
12     for all $\mathbbm{i}$, compute the loads excluding $u^{\mathrm{On}}_{\mathbbm{j}}$: $y^{-{\mathbbm{j}}}_{\mathbbm{i}}$, $\mathbbm{x}^{-{\mathbbm{j}}}_{\mathbbm{i}}$, and $n^{-{\mathbbm{j}}}_{\mathbbm{i}}$;
13
14    build the feasible action set $\mathcal{A}^{\mathrm{feas}}_{\mathbbm{j}}$ according to (23);
15
16    compute current utility $U_{\mathbbm{j}}(a^{\mathrm{cur}}_{\mathbbm{j}}\mid\bm{\pi}_{-\mathbbm{j}})$ via (19);
17
18    for each $a\in\mathcal{A}^{\mathrm{feas}}_{\mathbbm{j}}$ do
19        temporarily set $\pi_{\mathbbm{j}}\leftarrow a$ and evaluate $U_{\mathbbm{j}}(a\mid\bm{\pi}_{-\mathbbm{j}})$;
20
21
22    let $a^{\mathrm{best}}_{\mathbbm{j}}\in\arg\max_{a\in\mathcal{A}^{\mathrm{feas}}_{\mathbbm{j}}}U_{\mathbbm{j}}(a\mid\bm{\pi}_{-\mathbbm{j}})$;
23
24    if $U_{\mathbbm{j}}(a^{\mathrm{best}}_{\mathbbm{j}}\mid\bm{\pi}_{-\mathbbm{j}})-U_{\mathbbm{j}}(a^{\mathrm{cur}}_{\mathbbm{j}}\mid\bm{\pi}_{-\mathbbm{j}})>\varepsilon$ then
25        $\pi_{\mathbbm{j}}\leftarrow a^{\mathrm{best}}_{\mathbbm{j}}$;
26        $\mathrm{converged}\leftarrow\textbf{false}$;
27    else
28        $\pi_{\mathbbm{j}}\leftarrow a^{\mathrm{cur}}_{\mathbbm{j}}$;
29
30
31 $\mathbbm{t}\leftarrow\mathbbm{t}+1$;
32
Output: NE profile $\bm{\pi}^{\star}=\bm{\pi}$ and the associated allocation.
Algorithm 4 Proposed PG-BRD
Appendix E Proof Details for the Ordinal Potential and PG-BRD
This appendix collects the main technical ingredients behind the ordinal-potential formulation in the online stage and the design of PG-BRD. We have: (i) rewrite the model using a global assignment matrix, (ii) relate pointwise congestion penalties to integrated congestion via a simple sandwich inequality, (iii) characterize single-user deviations and the linearized soft-deadline utility, (iv) prove the ordinal-potential property and finite-improvement path (FIP), and (v) clarify why monetary payments are excluded from the potential.
E-A Matrix Formulation Based on the Global Assignment Matrix
For some of the theoretical arguments and for vectorized implementations, it is convenient to represent the joint offloading profile by a single global matrix.
Global assignment matrix. Define
$\bm{X}\in\{0,1\}^{\,|\mathcal{U}^{\mathrm{On}}|\times(1+|\mathcal{S}^{\mathrm{On}}|)}$ (51)
by stacking each SD’s binary decision vector $\bm{X}^{\mathrm{On}}_{\mathbbm{j}}=(x^{\mathrm{On}}_{\mathbbm{j}},x^{\mathrm{On}}_{\mathbbm{j},1},\ldots,x^{\mathrm{On}}_{\mathbbm{j},|\mathcal{S}^{\mathrm{On}}|})$. The $(\mathbbm{j},0)$-th entry corresponds to local execution ($\pi_{\mathbbm{j}}=0$), and the $(\mathbbm{j},\mathbbm{i})$-th entry corresponds to offloading to $s^{\mathrm{On}}_{\mathbbm{i}}$ ($\pi_{\mathbbm{j}}=\mathbbm{i}$). Under the one-hot per-row constraint $\sum_{\mathbbm{i}}x^{\mathrm{On}}_{\mathbbm{j},\mathbbm{i}}+x^{\mathrm{On}}_{\mathbbm{j}}=1$, $\bm{X}$ and $\bm{\pi}$ are equivalent representations.
Aggregates in matrix form. Using $\bm{X}$, the effective congestion load and physical workload at SP $s^{\mathrm{On}}_{\mathbbm{i}}$ are
$y_{\mathbbm{i}}(\bm{X})=\sum_{\mathbbm{j}\in\mathcal{U}^{\mathrm{On}}}\gamma_{\mathbbm{j},\mathbbm{i}}\,x^{\mathrm{On}}_{\mathbbm{j},\mathbbm{i}},\qquad\mathbbm{x}_{\mathbbm{i}}(\bm{X})=\sum_{\mathbbm{j}\in\mathcal{U}^{\mathrm{On}}}r_{\mathbbm{j}}\,x^{\mathrm{On}}_{\mathbbm{j},\mathbbm{i}}.$ (52)
These expressions are directly usable in optimization-based or vectorized implementations of PG-BRD.
Potential in matrix form. Substituting (52) into the potential in (21), we obtain the equivalent representation
$\Phi(\bm{X})$ $=\sum_{\mathbbm{j}\in\mathcal{U}^{\mathrm{On}}}\sum_{\mathbbm{i}\in\mathcal{S}^{\mathrm{On}}}S_{\mathbbm{j}}(\mathbbm{i})\,x^{\mathrm{On}}_{\mathbbm{j},\mathbbm{i}}-\sum_{\mathbbm{i}\in\mathcal{S}^{\mathrm{On}}}\int_{0}^{\,y_{\mathbbm{i}}(\bm{X})}\tau_{\mathbbm{i}}(z)\,dz$ (53)
$\quad-\sum_{\mathbbm{i}\in\mathcal{S}^{\mathrm{On}}}K_{\mathbbm{i}}\!\big(\mathbbm{x}_{\mathbbm{i}}(\bm{X})\big),$
which coincides with $\Phi(\bm{\pi})$ whenever each row of $\bm{X}$ has exactly one nonzero entry. In this form, $\Phi(\bm{X})$ decomposes into: (i) a linear term in non-congestion benefits, (ii) an integral term in congestion loads, and (iii) a convex term in SP costs.
E-B Integral-Pointwise Sandwich Inequality
To relate user-level congestion penalties to the integrated congestion term in the potential, we use a simple “sandwich” inequality for increasing functions.
Proposition 8 (Integral-pointwise sandwich).
Let $\tau_{\mathbbm{i}}:[0,\infty)\to[0,\infty)$ be nonnegative and increasing. For any $y\geq 0$ and any $\gamma_{\mathbbm{j},\mathbbm{i}}>0$,
$\gamma_{\mathbbm{j},\mathbbm{i}}\,\tau_{\mathbbm{i}}\!\big(y_{\mathbbm{i}}\big)\leq\int_{y_{\mathbbm{i}}}^{\,y_{\mathbbm{i}}+\gamma_{\mathbbm{j},\mathbbm{i}}}\tau_{\mathbbm{i}}(z)\,dz\leq\gamma_{\mathbbm{j},\mathbbm{i}}\,\tau_{\mathbbm{i}}\!\big(y_{\mathbbm{i}}+\gamma_{\mathbbm{j},\mathbbm{i}}\big).$ (54)
Sketch.
For all $z\in[y_{\mathbbm{i}},y_{\mathbbm{i}}+\gamma_{\mathbbm{j},\mathbbm{i}}]$, monotonicity gives $\tau_{\mathbbm{i}}(y_{\mathbbm{i}})\leq\tau_{\mathbbm{i}}(z)\leq\tau_{\mathbbm{i}}(y_{\mathbbm{i}}+\gamma_{\mathbbm{j},\mathbbm{i}})$. Integrating over this interval yields (54). Equivalently, by the mean-value theorem for integrals, the integral equals $\gamma_{\mathbbm{j},\mathbbm{i}}\tau_{\mathbbm{i}}(\xi)$ for some $\xi\in[y_{\mathbbm{i}},y_{\mathbbm{i}}+\gamma_{\mathbbm{j},\mathbbm{i}}]$, which directly implies the bounds. ∎
Interpretation. The integral in (54) can be viewed as the average of pointwise congestion penalties between $y_{\mathbbm{i}}$ and $y_{\mathbbm{i}}+\gamma_{\mathbbm{j},\mathbbm{i}}$. Its sign changes exactly when the pointwise terms do, which is the key to aligning changes in the potential with changes in individual utilities.
For the linear-response model $\tau_{\mathbbm{i}}(y)=y/f^{\mathrm{On}}_{\mathbbm{i}}$,
$\int_{y}^{\,y+\gamma}\!\!\tau_{\mathbbm{i}}(z)\,dz=\frac{(y+\gamma)^{2}-y^{2}}{2f^{\mathrm{On}}_{\mathbbm{i}}}=\frac{y\gamma}{f^{\mathrm{On}}_{\mathbbm{i}}}+\frac{\gamma^{2}}{2f^{\mathrm{On}}_{\mathbbm{i}}},$ (55)
so the integrated congestion level is $(y+\tfrac{\gamma}{2})/f^{\mathrm{On}}_{\mathbbm{i}}$, which lies between $\tau_{\mathbbm{i}}(y)$ and $\tau_{\mathbbm{i}}(y+\gamma)$.
E-C Single-User Deviation: Utility and Potential Changes
We next spell out the single-user deviation that underlies the ordinal-potential proof. The goal is to compare the sign of the utility change with that of the potential change.
Deviation setup. Fix an action profile $\bm{\pi}$ and SD $u^{\mathrm{On}}_{\mathbbm{j}}$. Consider a unilateral deviation
$\pi_{\mathbbm{j}}=\mathbbm{i}\ \longrightarrow\ \pi_{\mathbbm{j}}^{\prime}=\mathbbm{i}^{\prime},\qquad\bm{\pi}^{\prime}=(\bm{\pi}_{-\mathbbm{j}},\mathbbm{i}^{\prime}).$ (56)
By (20), the congestion loads at the two involved SPs satisfy
$y_{\mathbbm{i}^{\prime}}(\bm{\pi}^{\prime})$ $=y_{\mathbbm{i}^{\prime}}(\bm{\pi})+\gamma_{\mathbbm{j},\mathbbm{i}^{\prime}},$ (57)
$y_{\mathbbm{i}}(\bm{\pi}^{\prime})$ $=y_{\mathbbm{i}}(\bm{\pi})-\gamma_{\mathbbm{j},\mathbbm{i}}.$ (58)
The corresponding physical workloads evolve as
$x_{\mathbbm{i}^{\prime}}(\bm{\pi}^{\prime})=x_{\mathbbm{i}^{\prime}}(\bm{\pi})+r_{\mathbbm{j}},\qquad\mathbbm{x}_{\mathbbm{i}}(\bm{\pi}^{\prime})=\mathbbm{x}_{\mathbbm{i}}(\bm{\pi})-r_{\mathbbm{j}}.$ (59)
Utility change. Using the unified utility (19), the change in SD $u^{\mathrm{On}}_{\mathbbm{j}}$’s utility is
$\Delta U_{\mathbbm{j}}$ $=U_{\mathbbm{j},\mathbbm{i}^{\prime}}-U_{\mathbbm{j},\mathbbm{i}}$ (60)
$=\Big(S_{\mathbbm{j}}(\mathbbm{i}^{\prime})-S_{\mathbbm{j}}(\mathbbm{i})\Big)-\kappa_{\mathbbm{j}}\theta_{\mathbbm{j},\mathbbm{i}^{\prime}}\,\tau_{\mathbbm{i}^{\prime}}\!\big(y_{\mathbbm{i}^{\prime}}(\bm{\pi}^{\prime})\big)+\kappa_{\mathbbm{j}}\theta_{\mathbbm{j},\mathbbm{i}}\,\tau_{\mathbbm{i}}\!\big(y_{\mathbbm{i}}(\bm{\pi})\big)$
$\quad-\big(p_{\mathbbm{j},\mathbbm{i}^{\prime}}-p_{\mathbbm{j},\mathbbm{i}}\big).$
The first term is the change in non-congestion benefit, the next two terms capture the congestion effect, and the last term is a pure payment transfer.
Potential change. By (21), the potential change is
$\Delta\Phi$ $=\Phi(\bm{\pi}^{\prime})-\Phi(\bm{\pi})$
$=\underbrace{S_{\mathbbm{j}}(\mathbbm{i}^{\prime})-S_{\mathbbm{j}}(\mathbbm{i})}_{\text{non-congestion benefit}}$
$\quad-\underbrace{\Big(\int_{y_{\mathbbm{i}^{\prime}}(\bm{\pi})}^{\,y_{\mathbbm{i}^{\prime}}(\bm{\pi})+\gamma_{\mathbbm{j},\mathbbm{i}^{\prime}}}\!\!\tau_{\mathbbm{i}^{\prime}}(z)\,dz-\int_{y_{\mathbbm{i}}(\bm{\pi})-\gamma_{\mathbbm{j},\mathbbm{i}}}^{\,y_{\mathbbm{i}}(\bm{\pi})}\!\!\tau_{\mathbbm{i}}(z)\,dz\Big)}_{\text{integrated congestion variation}}$
$\quad-\underbrace{\Big(K_{\mathbbm{i}^{\prime}}\big(x_{\mathbbm{i}^{\prime}}(\bm{\pi})+r_{\mathbbm{j}}\big)-K_{\mathbbm{i}^{\prime}}\big(x_{\mathbbm{i}^{\prime}}(\bm{\pi})\big)\Big)}_{\text{SP $\mathbbm{i}^{\prime}$ cost increase}}$
$\quad+\underbrace{\Big(K_{\mathbbm{i}}\big(\mathbbm{x}_{\mathbbm{i}}(\bm{\pi})\big)-K_{\mathbbm{i}}\big(\mathbbm{x}_{\mathbbm{i}}(\bm{\pi})-r_{\mathbbm{j}}\big)\Big)}_{\text{SP $s^{\mathrm{On}}_{\mathbbm{i}}$ cost decrease}}.$ (61)
Payment terms are absent from $\Delta\Phi$ by construction (see Appx. E-I).
E-D Approximate Linearization of Soft-Deadline Utility
We briefly justify the linearized MU utility in (18) and the definition of the congestion-sensitivity coefficient $\kappa_{\mathbbm{j}}$.
Recall that for MU $u^{\mathrm{On}}_{\mathbbm{j}}$ offloading to $s^{\mathrm{On}}_{\mathbbm{i}}$,
$G^{\mathrm{M}}_{\mathbbm{j},\mathbbm{i}}=v_{\mathbbm{j}}\,\frac{\big[\tau^{\mathrm{ddl}}_{\mathbbm{j}}-T^{\mathrm{edge}}_{\mathbbm{j},\mathbbm{i}}\big]_{+}}{\tau^{\mathrm{ddl}}_{\mathbbm{j}}},\qquad T^{\mathrm{edge}}_{\mathbbm{j},\mathbbm{i}}=t^{\mathrm{comp}}_{\mathbbm{j},\mathbbm{i}}+t^{\mathrm{tx}}_{\mathbbm{j},\mathbbm{i}}.$ (62)
In the practically relevant regime where $T^{\mathrm{edge}}_{\mathbbm{j},\mathbbm{i}}\lesssim\tau^{\mathrm{ddl}}_{\mathbbm{j}}$, the operator $\cdot\,$$_{+}$ is active and
$G^{\mathrm{M}}_{\mathbbm{j},\mathbbm{i}}=v_{\mathbbm{j}}-v_{\mathbbm{j}}\frac{t^{\mathrm{tx}}_{\mathbbm{j},\mathbbm{i}}}{\tau^{\mathrm{ddl}}_{\mathbbm{j}}}-v_{\mathbbm{j}}\frac{t^{\mathrm{comp}}_{\mathbbm{j},\mathbbm{i}}}{\tau^{\mathrm{ddl}}_{\mathbbm{j}}}.$ (63)
Using $t^{\mathrm{comp}}_{\mathbbm{j},\mathbbm{i}}=\theta_{\mathbbm{j},\mathbbm{i}}\tau_{\mathbbm{i}}(y_{\mathbbm{i}})$, we obtain
$G^{\mathrm{M}}_{\mathbbm{j},\mathbbm{i}}=\underbrace{v_{\mathbbm{j}}-v_{\mathbbm{j}}\frac{t^{\mathrm{tx}}_{\mathbbm{j},\mathbbm{i}}}{\tau^{\mathrm{ddl}}_{\mathbbm{j}}}}_{\text{independent of }y_{\mathbbm{i}}}\;-\;\underbrace{v_{\mathbbm{j}}\frac{\theta_{\mathbbm{j},\mathbbm{i}}}{\tau^{\mathrm{ddl}}_{\mathbbm{j}}}\,\tau_{\mathbbm{i}}\big(y_{\mathbbm{i}}\big)}_{\text{congestion-dependent term}}.$ (64)
Adding the energy-saving term $c^{\mathrm{M},\mathrm{save}}_{\mathbbm{j},\mathbbm{i}}=E^{\mathrm{loc}}_{\mathbbm{j}}-E^{\mathrm{tx}}_{\mathbbm{j},\mathbbm{i}}$ with coefficient $\mathbb{V}^{\mathrm{M}}_{e}$ and subtracting the payment, the full MU utility can be written as
$U^{\mathrm{M}}_{\mathbbm{j},\mathbbm{i}}=\bar{S}_{\mathbbm{j}}(\mathbbm{i})\;-\;\kappa_{\mathbbm{j}}\,\theta_{\mathbbm{j},\mathbbm{i}}\,\tau_{\mathbbm{i}}\big(y_{\mathbbm{i}}\big)\;-\;p_{\mathbbm{j},\mathbbm{i}},$ (65)
where
$\bar{S}_{\mathbbm{j}}(\mathbbm{i})=v_{\mathbbm{j}}-v_{\mathbbm{j}}\frac{t^{\mathrm{tx}}_{\mathbbm{j},\mathbbm{i}}}{\tau^{\mathrm{ddl}}_{\mathbbm{j}}}+\mathbb{V}^{\mathrm{M}}_{e}\,c^{\mathrm{M},\mathrm{save}}_{\mathbbm{j},\mathbbm{i}},\qquad\kappa_{\mathbbm{j}}=\frac{v_{\mathbbm{j}}}{\tau^{\mathrm{ddl}}_{\mathbbm{j}}}>0.$ (66)
Thus, MUs share the same congestion structure as HUs, with user-specific $\kappa_{\mathbbm{j}}$. This linearization preserves the ordinal ranking of actions induced by the original soft-deadline utility in the operating regime of interest, and hence does not change the set of NE.
E-E Sign Consistency of Congestion Terms
The ordinal-potential property relies on the fact that congestion terms in the utility and in the potential move in the same direction.
Proposition 9 (Sign consistency).
For deviation $\pi_{\mathbbm{j}}:\mathbbm{i}\to\mathbbm{i}^{\prime}$, the congestion part of the utility change,
$\kappa_{\mathbbm{j}}\theta_{\mathbbm{j},\mathbbm{i}^{\prime}}\,\tau_{\mathbbm{i}^{\prime}}\!\big(y_{\mathbbm{i}^{\prime}}(\bm{\pi}^{\prime})\big)+\kappa_{\mathbbm{j}}\theta_{\mathbbm{j},\mathbbm{i}}\,\tau_{\mathbbm{i}}\!\big(y_{\mathbbm{i}}(\bm{\pi})\big),$
and the congestion-integral term in the potential change,
$\Big(\!\int_{y_{\mathbbm{i}^{\prime}}(\bm{\pi})}^{\,y_{\mathbbm{i}^{\prime}}(\bm{\pi})+\gamma_{\mathbbm{j},\mathbbm{i}^{\prime}}}\!\!\tau_{\mathbbm{i}^{\prime}}(z)\,dz-\int_{y_{\mathbbm{i}}(\bm{\pi})-\gamma_{\mathbbm{j},\mathbbm{i}}}^{\,y_{\mathbbm{i}}(\bm{\pi})}\!\!\tau_{\mathbbm{i}}(z)\,dz\!\Big),$
have the same sign.
Proof.
Apply Proposition 8 to both integrals. Each integral lies between a lower and an upper pointwise congestion term. After scaling by the positive factors $\kappa_{\mathbbm{j}}\theta_{\mathbbm{j},\mathbbm{i}}/\gamma_{\mathbbm{j},\mathbbm{i}}$ and $\kappa_{\mathbbm{j}}\theta_{\mathbbm{j},\mathbbm{i}^{\prime}}/\gamma_{\mathbbm{j},\mathbbm{i}^{\prime}}$, the corresponding bounds match the pointwise expressions in the lemma. Hence the two congestion components must share the same sign. ∎
For the linear model $\tau_{\mathbbm{i}}(y)=y/f^{\mathrm{On}}_{\mathbbm{i}}$, the result is even more transparent: the integral is simply the area under a straight line, so shifting traffic to a more congested SP increases both the user’s delay and the integrated congestion cost.
E-F Monotonicity and Convexity of SP Cost
The SP-cost term $K_{\mathbbm{i}}(\cdot)$ is chosen to discourage load concentration on heavily utilized SPs.
Assume $K_{\mathbbm{i}}:\mathbb{R}_{+}\to\mathbb{R}_{+}$ is increasing and convex. Then for $x^{\prime}_{\mathbbm{i}}=\mathbbm{x}_{\mathbbm{i}}-r_{\mathbbm{j}}$ and $x^{\prime}_{\mathbbm{i}^{\prime}}=x_{\mathbbm{i}^{\prime}}+r_{\mathbbm{j}}$,
$K_{\mathbbm{i}^{\prime}}(x_{\mathbbm{i}^{\prime}}+r_{\mathbbm{j}})-K_{\mathbbm{i}^{\prime}}(x_{\mathbbm{i}^{\prime}})\geq 0,\qquad K_{\mathbbm{i}}(\mathbbm{x}_{\mathbbm{i}})-K_{\mathbbm{i}}(\mathbbm{x}_{\mathbbm{i}}-r_{\mathbbm{j}})\geq 0.$ (67)
Convexity further implies that the marginal cost $K_{\mathbbm{i}}(x+\Delta)-K_{\mathbbm{i}}(x)$ is nondecreasing in $x$: adding the same workload to a more congested SP incurs a larger cost. Consequently,
$\Big[K_{\mathbbm{i}^{\prime}}(x_{\mathbbm{i}^{\prime}}+r_{\mathbbm{j}})-K_{\mathbbm{i}^{\prime}}(x_{\mathbbm{i}^{\prime}})\Big]+\Big[K_{\mathbbm{i}}(\mathbbm{x}_{\mathbbm{i}})-K_{\mathbbm{i}}(\mathbbm{x}_{\mathbbm{i}}-r_{\mathbbm{j}})\Big]$ (68)
moves in a direction consistent with the deviation: shifting load towards a more congested SP is penalized in the potential.
For the linear model $K_{\mathbbm{i}}(x)=\omega_{1}\,e^{\mathrm{comp}}_{\mathbbm{i}}\,x/f^{\mathrm{On}}_{\mathbbm{i}}$, these increments are simply proportional to $r_{\mathbbm{j}}$ and preserve the sign.
E-G Proof Sketch of Proposition 7
We are now ready to justify the ordinal-potential property.
Consider a deviation $\pi_{\mathbbm{j}}:\mathbbm{i}\to\mathbbm{i}^{\prime}$ that strictly increases SD $u^{\mathrm{On}}_{\mathbbm{j}}$’s utility, i.e., $\Delta U_{\mathbbm{j}}>0$ in (60). Comparing (60) and (61), we note:
  • The non-congestion term $S_{\mathbbm{j}}(\mathbbm{i}^{\prime})-S_{\mathbbm{j}}(\mathbbm{i})$ appears identically in both $\Delta U_{\mathbbm{j}}$ and $\Delta\Phi$.
  • By Proposition 9, the congestion contributions to $\Delta U_{\mathbbm{j}}$ and $\Delta\Phi$ have the same sign.
  • By Appx. E-F, the SP-cost variation responds monotonically and convexly to the direction of load shift, further aligning the sign of $\Delta\Phi$ with that of $\Delta U_{\mathbbm{j}}$.
  • Payment differences $(p_{\mathbbm{j},\mathbbm{i}^{\prime}}-p_{\mathbbm{j},\mathbbm{i}})$ only redistribute welfare and are excluded from $\Delta\Phi$ (Appx. E-I), so they do not affect the ordinal comparison.
Hence, $\Delta U_{\mathbbm{j}}>0$ implies $\Delta\Phi>0$, which proves that $\Phi(\bm{\pi})$ is an ordinal potential for the considered game, as stated in Proposition 7.
E-H Finite-Improvement Property and Existence of NE
The ordinal-potential property directly yields convergence of better-response dynamics and existence of pure-strategy Nash equilibria.
Let $\mathcal{A}_{\mathbbm{j}}=\{0\}\cup\mathcal{S}^{\mathrm{On}}$ be SD $u^{\mathrm{On}}_{\mathbbm{j}}$’s finite action set, and $\mathcal{A}=\prod_{\mathbbm{j}}\mathcal{A}_{\mathbbm{j}}$ the joint action space. Consider any strict better-response dynamics in which, at each step, some SD switches to an action that strictly improves its utility. By Proposition 7, every such deviation strictly increases the potential $\Phi(\bm{\pi})$.
Since $\mathcal{A}$ is finite, the sequence of potential values cannot increase indefinitely. The process must terminate after finitely many steps at a profile $\bm{\pi}^{\star}$ where no SD can further increase its utility via a unilateral deviation. By definition, $\bm{\pi}^{\star}$ is a pure-strategy NE. Moreover, any local maximizer of $\Phi(\bm{\pi})$ is a pure-strategy NE, because any unilateral move would reduce (or not increase) $\Phi$, and hence would not increase the deviator’s utility in an ordinal-PG.
This finite-improvement property underpins PG-BRD: as long as SDs occasionally update their actions based on best responses, the algorithm converges in finitely many steps to a stable congestion-aware offloading configuration.
E-I Treatment of Monetary Payments in the Potential Function
We finally clarify why monetary payments are left out of the potential, which is sometimes questioned from a mechanism-design perspective.
Let $\mathcal{W}(\bm{\pi})$ denote the non-transfer welfare:
$\mathcal{W}(\bm{\pi})=\sum_{\mathbbm{j}:\,\pi_{\mathbbm{j}}\neq 0}S_{\mathbbm{j}}(\pi_{\mathbbm{j}})-\sum_{\mathbbm{i}}\int_{0}^{\,y_{\mathbbm{i}}(\bm{\pi})}\!\!\tau_{\mathbbm{i}}(z)\,dz-\sum_{\mathbbm{i}}K_{\mathbbm{i}}\!\big(\mathbbm{x}_{\mathbbm{i}}(\bm{\pi})\big).$ (69)
Agent-to-agent payments $\{p_{\mathbbm{j},\mathbbm{i}}\}$ are internal transfers between SDs and SPs. In the absence of external taxes or subsidies, they sum to zero at the system level and do not affect $\mathcal{W}(\bm{\pi})$. Thus, payments can be omitted from the potential without changing ordinal comparisons between profiles.
In other words, $\Phi(\bm{\pi})$ is designed as a proxy for real system efficiency, capturing latency savings, energy savings, congestion, and SP costs, while treating payments as a separate distributional layer superimposed on top of this efficiency metric. This follows standard practice in congestion and resource-allocation games and allows PG-BRD to focus on aligning individual incentives with global performance, independently of how the resulting welfare is split among participants.